### v2

# Instruction Fine-Tuning: Does Prompt Loss Matter?

Abstract
We present a study analyzing the effects of prompt loss weighting (PLW) on supervised instruction fine-tuning. We recreated Stanford’s Alpaca experiment with both LLaMA 1 and LLaMA 2 and multiple instruction datasets. We found that performance of models fine-tuned on our short-completion dataset had a statistically-significant negative quadratic relationship with PLW, but performance of models fine-tuned on medium- and long-completion data did not show any relationship with PLW. I.e., prompt loss can be safely ignored for many datasets. For short-completion data, small values ($0.01-0.1$) of PLW were optimal for multiple-choice and short-generation tasks while large values ($\approx 1.0$) of PLW were optimal for long-generation tasks. We concluded that low non-zero PLW encourages models to not diverge from pre-trained model weights during training and high PLW reduces overfitting. Finally, we present a rough guide for selecting PLW values based on the completion-prompt length ratio of fine-tuning data.
Instruction Fine-Tuning: Does Prompt Loss Matter?
Mathew Huerta-Enochian EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul mathew@eq4all.co.kr
1 Introduction
Recent research in language modeling has made huge advances in training instruction-following agents. Both supervised fine-tuning and reinforcement learning have been employed to much success. However, our understanding of optimal hyperparameters and standards of practice have been slow to catch up. This research intends to contribute to hyperparameter selection for supervised instruction fine-tuning (SIFT) via an in-depth analysis of a single training hyperparameter: prompt loss weight (PLW).
The goal of SIFT is to teach a model to perform a language task given a prompted instruction. While training, model parameters are updated by optimizing for next-token maximal likelihood classification. However, it is unclear if this optimization should be performed on the entire prompt-completion sequence or only on the target completion.
Most open sourced solutions for supervised fine-tuning (SFT) either mask the prompt loss or use the entire sequence loss. On the other hand, OpenAI previously supported a prompt_loss_weight parameter that allowed users to specify PLW during SIFT.
We make the following contributions related to this parameter:
  • We show that PLW has a statistically significant negative quadratic relationship with model performance when fine-tuning on short-completion instruction data.
  • We show that PLW and prompt masking parameters can be safely ignored when fine-tuning on medium- or long-completion data, streamlining the fine-tuning process.
  • We provide evidence that the benefits of PLW for short-completion data are due to a regularizing effect and not due to previously suspected mechanisms.
  • We present a baseline for PLW selection for short-completion fine-tuning.
We provide relevant background and hypotheses in section 2 and section 3, respectively. We resent our methodology in section 4. Descriptive and regression analysis and an exploration of causal mechanisms are presented in section 5. We extend the experiment to additional datasets and present a basic model for prompt loss weight selection in section 6. We present conclusions in section 7.
2 Background
2.1 Definitions
We define instruction data as one or many instances of structured text data, each containing an instruction text, an optional context or input text, and a target output text. For the rest of the paper, we use the term prompt to refer to the concatenation of instruction and input texts and the term completion to refer to the target output text. SIFT attempts to train a model to generate an appropriate completion for a given prompt.
We define the generation ratio $R_{g}$ as the ratio of completion length to prompt length. We then divide instruction data into two broad categories. Data with $R_{g}<1$ are short-completion data, and data with $R_{g}\geq 1$ are long-completion data. When applied to an entire dataset, we take $R_{g}$ to be the mean completion-prompt ratio.
2.2 Relevant Research and Libraries
HuggingFace’s Transformers library Wolf et al. (2020), the de facto library for training LLMs, allows users to mask select tokens when calculating loss. In Transformers, optimizing next-token prediction on prompt tokens is therefore binary—either token loss is masked (PLW $=0$) or it is unmasked (PLW $=1$).
Until recently, OpenAI supported a prompt_loss_weight parameter in their fine-tuning API, but it was officially removed as part of the v1 fine_tune API deprecation in early January, 2024. This prompt_loss_weight parameter used a default value of $0.01$ with the following parameter explanation:
“This controls how much the model tries to learn to generate the prompt (as compared to the completion which always has a weight of 1.0), and can add a stabilizing effect to training when completions are short. If prompts are extremely long (relative to completions), it may make sense to reduce this weight so as to avoid over-prioritizing learning the prompt.”
Though we could not find a study validating OpenAI’s claim or any literature that presents an analysis of PLW, there were several studies that reported parameter values for prompt_loss_weight when fine-tuning OpenAI models.
Though they do not provide their reasoning, Kozachek (2023) reported that they fine-tuned GPT-3 with a prompt_loss_weight of $0.1$. Dodgson et al. (2023) reported using the default value of $0.01$ when fine-tuning GPT models. Wang et al. (2023b) reported that a PLW of 00 performed best for them when working on the Self-Instruct framework. Interestingly, Wutschitz et al. (2023) reported hyperparameter search results for next-sentence-prediction on Elsevier data using PLWs of $0.1$ and $0.5$ and found $0.5$ to give the best results. Similar to OpenAI’s deprecated API, BLoomAI’s API supports a prompt_loss_weight parameter with a default value of 0.01.
Mean (Std) Tokens
Dataset Instruction Input Completion Total Tokens $R_{g}$
AlpacaData 13.40 (4.97) 6.25 (14.38) 64.51 (64.85) 4,376,482 3.27
AlpacaDataCleaned 14.21 (9.95) 6.42 (17.65) 162.74 (150.89) 9,490,837 7.83
AlpacaDataShort 16.93 (13.10) 162.34 (151.69) 14.62 (10.99) 10,035,667 0.082
Table 1: Dataset statistics: mean tokenized instruction, input, and completion sequence lengths (standard deviations in parentheses), total token counts for each dataset, and the generation ratio $R_{g}$.
3 Hypotheses
Based on OpenAI’s documentation for prompt_loss_weight, we expect PLW to have a stabilizing effect on LLM SIFT when using short-completion data. Our expectation was that PLW preserves the hidden representation learned for next-token-classification when most embeddings are being used to predict tokens at the end of the sequence (as is the case when fine-tuning on short-completion data). This would represent a positive relationship between PLW and downstream performance.
However, unless the model is expected to generate prompt-like data in downstream tasks, the training and downstream sequence contexts will not match. In other words, training a model to maximize next-token-prediction on prompt tokens should be most useful for generating instruction data. Due to this train-test mismatch, we expect PLW to also have a negative influence on downstream performance.
Based on the assumption that some amount of PLW is beneficial for short-completion SIFT, we expect these two competing factors to result in a downward curved relationship between downstream performance and PLW. Limiting PLW to the range of [0, 1], we postulate that there is a critical value $\lambda$ for PLW with $10<=\lambda<=$ 1. For PLW less than $\lambda$, the positive effect dominates the negative effect and for values greater than $\lambda$, the negative effect dominates the positive effect. If $\lambda=0$, then PLW’s contribution to model performance is strictly negative. Conversely, if $\lambda=1$, then PLW contributes strictly positively to model performance.
We postulate that both $\lambda$ and the magnitude of the effect are inversely proportional to $R_{g}$ since there will be $R_{g}$ as many completion token loss terms than prompt loss terms. Note that $\lambda$ is not an intrinsic characteristic of the dataset, model architecture, or training algorithm. Rather, it would depend on numerous factors and change for each task.
We tested the relationship between PLW and model performance using three fine-tuning datasets spanning a range of $R_{g}$ values: 0.327, 7.83, and 0.082. We present the following hypotheses to be tested against each dataset.
Null Hypothesis ($\mathbf{H_{0}}$) Prompt loss weight has no relationship with model performance.
Alternative ($\mathbf{H_{1}}$) Prompt loss weight has a quadratic relationship with model performance.
We fixed the threshold $\alpha=0.05$ significance level, and we expected to be able to reject $\mathbf{H_{0}}$ for models trained on short-completion data.
4 Methodology
To evaluate the effect of PLW on downstream performance, we used a factorial design methodology and repeated the Alpaca experiment Taori et al. (2023) with three experimental variables. We tested ten discrete levels of PLW, two pre-trained language models (PTLMs), and three instruction fine-tuning datasets for a total of sixty experimental training runs and evaluated each run on thirteen benchmarks.
We used the original Alpaca code and Transformers library, only modified to add PLW. Training was performed exactly as per the original Alpaca experiment, and we used the hyperparameters suggested by the authors, modifying only the three experimental parameters (PLW, PTLM, dataset) with each run.
We provide additional details for reproducibility in appendix C and will release our trained models on the HuggingFace Hub.
4.1 Prompt Loss Weight
We limited our evaluation of PLW to factors in the range [0, 1], focusing on values close to zero:
[…]
$1{\times}10^{-2},2.463{\times}10^{-2},5{\times}10^{-2},$
$1{\times}10^{-1},2.463{\times}10^{-1},5{\times}10^{-1},1.0\}$
Note that PLW $=0.0$ is identical to the masking used in the original Alpaca project, and PLW $=1.0$ is equivalent to unmasked training.
For all analysis, we transformed our PLW values to be closer to uniform on the interval [0, 1] using a power function
[…]
where the power $p=0.30103$ was chosen semi-arbitrarily such that $f(0.1)=0.5$. We denote the transformed PLW values as wp
4.2 Pre-Trained Language Model
We fine-tune both LLaMA 1 7B Touvron et al. (2023a) to recreate the original Alpaca experiment and LLaMA 2 7B Touvron et al. (2023b) to provide more relevant results.
4.3 Fine-Tuning Dataset
We ran all experiments with three datasets: AlpacaData (the instruction dataset from the original Alpaca experiment), AlpacaDataCleaned, and AlpacaDataShort.
AlpacaDataCleaned Ruebsamen (2023) is a cleaned and curated version of AlpacaData that has recently been combined with data from the GPT4 LLM dataset Peng et al. (2023).
Cleaning is noted as ongoing and includes fixes for the following issues in AlpacaData: hallucinations, merged instructions, empty outputs, empty code examples, instructions to generate images, N/A outputs, inconsistent input fields, wrong answers, nonsensical instructions, and extraneous control characters.
We generated AlpacaDataShort from AlpacaDataCleaned by swapping instructions and completions for all long-completion instances and rephrasing the task as an instruction-prediction task.
For instances with empty input, we used
Predict the prompt that generated the following AI output.
as the new instruction field. For instances with non-empty input, we used
Predict the prompt that generated the below AI output given the following context. Context: <original_input>
as the new instruction field. In all cases we then used the original completion as the new input field and the original instruction as the new completion field.
Descriptive statistics for tokenized instruction, input, and completion sequences are presented in table 1. Note that AlpacaDataCleaned is strongly long-completion with an $R_{g}$ of $7.83$ while AlpacaDataShort is short-completion with an $R_{g}$ of $0.082$.
4.4 Performance Evaluation
Task V. Shots Split Type
ARC Challenge* 0 25 Test MC
PIQA 0 0 Val MC
TruthfulQA-MC2* 1 6† Val MC
WinoGrande* 0 5 Val MC
TruthfulQA-Gen 1 6† Val GS
[…]
Alpaca Eval (Mixtral) 1 1 Test GL
PandaLM 1 0 Test GL
Table 2: Benchmark tasks used for evaluation. All benchmarks were scored using Eleuther AI’s Language Model Evaluation Harness, the AlpacaEval 1 framework, and the PandaLM framework. validation splits were used when test splits were unavailable, and validation and test splits were combined for noisy benchmarks. Benchmark completion type is noted here as “MC” for multiple choice, “GS” for short generation, and “GL” for long-generation.
*Task used in HuggingFace’s Open LLM Leaderboard.
†This benchmark was calculated with num_fewshot $=0$ but uses a built-in minimum of 6 shots.
We evaluated each model on thirteen instruction benchmarks covering both multiple choice and text generation tasks. We attempted to select benchmarks that were relatively cheap to compute and covered a range of tasks. We used three evaluation frameworks: EleutherAI’s Language Model Evaluation Harness (EEH) Gao et al. (2023), Alpaca Eval 1 Li et al. (2023), and PandaLM Wang et al. (2023a, 2024). See table 2 for details on benchmark tasks.
Eleven benchmarks were run using EEH. Four benchmarks weremultiple choice tasks: ARC Challenge Clark et al. (2018), PIQA Bisk et al. (2020), TruthfulQA-MC2 Lin et al. (2022), and WinoGrande Sakaguchi et al. (2019). Seven benchmarks were short generation tasks: TruthfulQA-Gen Lin et al. (2022) and six WMT14 and WMT16 translation benchmarks Bojar et al. (2014, 2016), limited to four languages the PTLMs learned during pretraining (English, French, German, and Romanian). For WMT benchmarks, we used the instruction “Translate the following from <src_lang> to <tgt_lang>” and evaluated over both the validation and test sets to reduce variance.
Long-generation performance was evaluated using Alpaca Eval 1 and PandaLM, which are both LLM-as-a-judge frameworks.
The default auto-evaluator for Alpaca Eval 1 is GPT4, but evaluation of all training runs would be beyond the scope of this research, so we used Mixtral 8X7B Jiang et al. (2024) as an auto-evaluator. Mixtral performed the best of all available LLMs that we tested on Alpaca Eval’s evaluator test dataset Dubois et al. (2023), scoring 64.9 for agreement with human evaluators, just below Claude’s Anthropic (2023) 65.3 agreement. For reference, the default GPT4 auto-evaluator scored a 70.99 human agreement.
5 Results and Discussion
(a) Simple Aggregate
(b) AlpacaData Relative Aggregate
(c) AlpacaDataClean Relative Aggregate
(d) AlpacaDataShort Relative Aggregate
Figure 1: Performance by transformed PLW. (a) A simple performance aggregate score (the unweighted mean of benchmark scores). (b), (c), (d) Relative aggregate performance scores where scores per task for each task and group are min-max scaled to show common trends, regardless of scale. Note that aggregate scores for only the AlpacaDataShort models show a relationship with transformed PLW. Best viewed in color.
5.1 Performance Trends
A visualization of the simple and relative performance aggregates by wp and dataset is presented in figure 1. The simple aggregate is the unweighted mean of all task scores for given wp and dataset values. For the relative aggregate, we first min-max normalized benchmark scores over each dataset-benchmark group and then performed aggregation. The simple aggregate is dominated by large performance changes on the AlpacaEval 1 and PandaLM benchmarks. Therefore, all analysis was performed using the relative aggregate and the simple aggregate was visualized only for completeness.
[…]
Identification of trends on the six translation benchmarks was less clear due to the high level of noise relative to performance range. Aggregating scores from benchmarks translating from English to the paired language benchmarks is suggestive of an increasing relationship, but is incredibly noisy.
The performance difference across wp values for the two long-generation benchmarks was around twenty percentage points, in stark contrast to the less than two percentage point change for short-generation and multiple choice benchmarks. This suggests that PLW plays an important role in the ability to generate high quality text, and the optimal PLW for short-generation and long-generation benchmarks is clearly different.
Finally, performance of LLaMA 2 models was higher than that of LLaMA 1 models and performance of AlpacaDataCleaned models were higher than that of AlpacaData models, validating the improvements of LLaMA 2 and AlpacaDataCleaned over their predecessors.
See Appendix A for detailed figures for each benchmark.
5.2 Regression
For each data group, we fit a generalized linear mixed model (GLMM) predicting min-max transformed benchmark scores. As with the min-max transformation for the relative aggregate, benchmark scores were min-max transformed over dataset-benchmark groups.
We expected a quadratic relationship between the score and wp, so we included a second order polynomial of wp as a fixed effect. Furthermore, we knew that the performance-PLW relationship varies by benchmark and that scores were min-max normalized over each benchmark, so we used a random slope with respect to benchmark. Since we did not min-max normalize over PTLM groups and since we saw consistent improvement when using LLaMA 2, we modeled a random intercept for PTLM. This resulted in the following equation that we fit with the R library glmmTMB:
[…]
Coeff
P-Value wp wp2 (Int)
AlpacaData 0.237 1.185 -0.917 (-0.131)
AlpacaDataCleaned 0.0861 1.238 -0.812 (-0.231)
AlpacaDataShort <0.001 5.590 -4.284 (-1.043)
Table 3: wp p-values and coefficients by training dataset. Statistically significant results are in bold. Note that though convergence warnings were raised for regression on both AlpacaData and AlpacaDataCleaned, coefficient and p-value scores are reported for completeness.
(a) Training Loss Stability
(b) Weight Distance
(c) Train Data Memorization
(d) AE Generation Length
Figure 2: Analysis of causal mechanism. Boxplots use the 0.25, 0.5, and 0.75 quantiles with whiskers at 0.09 and 0.91 quantiles. Best viewed in color. (a) Training Loss Stability: Relative Standard Deviation (RSD) of five-step training loss windows show increase instability for small (non-zero) PLWs. (b) Weight Distance: Distance between learned weights and PTLM weights is smaller for small (non-zero) PLWs. (c) Train Data Memorization: Completion Sacre BLEU scores on training data prompts as an indicator for overfitting. (d) AE Generation Length: Generation lengths on the Alpaca Eval test set for varying PLW values.
P-values and coefficients are presented in table 3. Regression on both AlpacaData and AlpacaDataCleaned produced convergence warnings and appropriate models could not be adequately fit. We tried reducing the complexity of the model, but no significant relationship with wp could be found. However, the model fit on AlpacaDataShort converged and passed residual normallity, homoscedasticity, and other checks for soundness. For the AlpacaDataShort case, min-max transformed performance showed a statistically significant negative quadratic relationship with wp at our target $\alpha=0.05$ significance level.
This means that while wee could not reject the null hypothesis for the AlpacaData and AlpacaDataCleaned scenarios, for the AlpacaDataShort scenario, there was sufficient evidence to reject the null hypothesis in favor of the alternative hypothesis H1.
Using the fixed effect coefficients, we can predict the critical PLW value $\lambda$ for AlpacaDataShort fine-tuning that maximizes the min-max transformed benchmark scores. The coefficients for wp2, wp, and the intercept were -4.28, 5.590, -1.043, respectively. We can rewrite this relationship as:
[…]
5.3 Causal Mechanism & Interpretation
To identify possible causal mechanisms, we first investigated the effects of PLW on training stability by analyzing training loss relative standard deviation (RSD) over all five-step windows. See figure 1(a) for a boxplot of mean RSD for each model. For all dataset and PTLM factors, increasing PLW from zero led to a sharp increase in mean RSD and then a slow decrease to a minimum mean RSD at PLW $=1$. There is no obvious explanation for why training loss RSD would initially increase before decreasing.
If training loss stability was the primary factor in improved performance, we would expect stability to be best for PLW between $0.01$ and $0.1$ and for performance at PLW $=0$ to be similar with performance at PLW $=0.01$ since mean RSD at these values are similar. However, mean RSD drops by a factor of two across the PLW $\in[0.01,0.1]$ range, and performance at PLW $=0.01$ is significantly higher than the masked prompt loss scenario. Training loss mean RSD is lowest at PLW $=1$, but performance on ARC Challenge, PIQA, WinoGrande, and TruthfulQA-Gen show clear decreasing trends at this value. Furthermore, the three tasks showing positive trends at PLW $=1$ cannot be adequately explained by this factor since performance increases regardless of loss stability.
There is likely either a tradeoff between training loss stability and some other factors that affects model performance or model loss stability is not an important factor. Considering that both AlpacaData and AlpacaDataCleaned models also showed a negative quadratic trend for mean loss RSD, we tentatively concluded that loss stability is not the driving factor for the modeled relationship.
We then checked if PLW was providing regularization to the weight update step, possibly improving performance by keeping weights close to the PTLM. See figure 1(b) for a visualization of weight distance from PTLM. Interestingly, for AlpacaDataShort, fine-tuned weights were closer to those of the PTLM for small values of PLW but were much farther for PLW $<0.0005$ and PLW $>0.1$. We would expect weights to change more when loss is erratic, but the range of PLW values that better preserved PTLM weights was similar to the range of increased training loss RSD. Unfortunately, we do not have an explanation for this apparent contradiction and simply conclude that PTLM model weights were better preserved for small non-zero PLW despite high loss instability.
We next explored how PLW affected training data memorization. We sampled 10,000 unique prompts from the AlpacaDataShort training set, generated completions from each prompt, and calculated corpus BLEU-4 scores. We found that for PLW from 0.0 to around 0.1, models memorized most of the training data, consistently scoring near 80 corpus BLEU. Corpus BLEU then decreased as PLW increased from 0.1. We also analyzed generation length on the AE1 test set which showed high deviations in generation length for the high loss RSD PLW values for LLaMA 2 only and a generally increasing length trend with PLW. Considering that AlpacaDataShort contains short completions, this strongly suggests that PLW contributes a regularizing effect and helps decrease overfitting.
The positive relationship between performance and PLW on Alpaca Eval and PandaLM as opposed to the negative quadratic relationship seen on several of the other benchmarks can then be explained by a tradeoff between longer generations and decreased overfitting for high PLW and better preservation of PTLM weights for low PLW.
6 Predicting Prompt Loss Weight
Figure 3: Ratio-PLW Relationship
We have shown that PLW is an important hyperparameter when fine-tuning on short-completion data but is effectively irrelevant when using long-completion data. We now present a model for predicting an optimal PLW given a dataset’s $R_{g}$.
We first extended the previous experiments to two additional datasets: AlpacaDataMedium and AlpacaDataTiny to increase coverage of the parameter space. AlpacaDataMedium and AlpacaDataTiny have $R_{g}$ values of 1.0 and 0.042, respectively, and were generated similarly with AlpacaDataShort, but with different target completion-prompt ratios.
We then fit several generalized additive model (GAMs) with a tensor smooth for the interaction between PLW and $R_{g}$. GAMs offer more flexible modeling than linear models but at the expense of interpretability. We fit our models using the R library mgcv and the following equation:
score ~ te(w, r, k=3) + factor(b)”, where te is a full tensor product smooth, w is the untransformed PLW parameter, r is the $R_{g}$, k is the number of splines, and b is the benchmark task.
Using the fitted w-r interaction, we can now estimate an optimal PLW value for a given completion-prompt ratio $R_{g}$. See figure 3 for a visualization.
Roughly, the fitted interaction term recommends using PLW $=0.155$ for small completion-prompt length ratios ($R_{g}\leq 1$) and up to PLW $=0.278$ for a $R_{g}=1.5$ for optimal performance across all tasks. This prediction is close to our regression-predicted value of 0.242. The interaction term also confirms our conclusion that PLW is less important for data with relatively long completions.
This is a rough guide, and we recommend using it as a starting point. See appendix B for additional plots using different benchmark tasks.
7 Conclusion
In this study, we explored the downstream effects of prompt loss weight (PLW) for LLM supervised instruction fine-tuning (SIFT).
We showed that performance is unaffected for the two long-completion datasets (AlpacaData, the original Alpaca dataset, and AlpacaDataCleaned, an updated version of AlpacaData) while performance had a negative quadratic relationship with prompt loss weight on our short-completion dataset (with a completion-prompt ratio $R_{g}\approx 0.08$).
Interestingly, the highest scores on the ARC Challenge, PIQA, and WinoGrande benchmarks were all from models trained on short-completion dataset. This suggests that given a good choice of prompt loss weight, long-completion training data may not be necessary for instruction fine-tuning LLMs for language understanding and short generation tasks. Conversely, using long-completion data should be more robust as prompt loss weight and masking parameters can be ignored.
We further analyzed the AlpacaDataShort models and found that low non-zero PLW discourages deviation from the PTLM weights and high PLW reduces overfitting and encourages longer generation. We suggest that these two factors are responsible for the negative quadratic relationship between performance and PLW.
Finally, we estimated optimal PLWs over a range of completion-prompt length ratios.
Limitations
  1. 1.
    We analyzed prompt loss weighting (PLW) for instruction fine-tuning LLMs. We characterized three fine-tuning datasets by their relative completion-prompt length ratios and reported on the effect of PLW when training on each dataset. It would be helpful to extend this research to a wider range of datasets to increase the strength of our conclusions and create more complete guidelines for prompt loss weighting.
  2. 2.
    We used a fixed seed $=42$ for all experiments. Since we used pre-trained models and no layers were freshly initialized, there was little variance in initial experiments. Therefore, we decided to limit runs to a single seed.
  3. 3.
    Suggested values for PLW from section 6 are based on the included experiments. Best PLW values when fine-tuning different models or using different datasets or training regimes may vary from the relationships shown here, though we are still confident that performance will not vary significantly by PLW for long-completion data.
  4. 4.
    LLM-as-evaluator approaches like PandaLM and Alpaca Eval 1 are still relatively new, and these approaches are still being actively developed. We chose to use Mixtral as an auto-evaluator for Alpaca Eval. Although this is not the recommended auto-evaluator, we felt its performance on the human-labeled test set was good enough to justify its use.
  5. 5.
    A more confident conclusion about the causative mechanism could not be made. As with most techniques in machine learning, the suggested causative relationship should be validated through numerous experiments and projects. Unfortunately, since we expect most instruction fine-tuning datasets to be long-completion datasets, further research into PLW will likely be limited.
Ethical Considerations
In this paper, we presented an analysis of the prompt loss weight hyperparameter for supervised instruction fine-tuning. We did not rely on human evaluators, and at no point in our research did we expose anyone to risk of harm.
We do acknowledge that standard deep learning training methods have a high carbon footprint, and we performed a total of 100 fine-tuning training runs. Model outputs cannot be predicted in advance, and, while we release our model weights in the spirit of transparency and collaboration, models may hallucinate or produce offensive output. Our synthetic dataset was generated from another publicly released dataset and we did not perform additional content filtering. A warning about both of these issues will be released along with the model and dataset.
Acknowledgements
This research was supported by Gwangju City’s Artificial Intelligence Industry Cluster Agency (AICA) and by the Korean Ministry of Trade, Energy, and Industry (MOTIE) under the KoSIGN project [Grant Number: 20014406].
This research was possible thanks to the support of my coworkers at EQ4ALL.
Appendix A Benchmark Performance Figures
This appendix contains individual benchmark performance plots and additional discussion of performance trends and results.
Overall, models trained on AlpacaDataCleaned performed best on the selected benchmarks, with models trained on AlpacaDataShort meeting or outperforming these models on numerous benchmarks for optimal selection of PLW. The general PTLM trend was for LLaMA 2 models to outperform LLaMA 1 models, again with some exceptions. In general, models trained on AlpacaData and AlpacaDataCleaned did not show consistent relationships with PLW on the tested benchmark.
The rest of this section focuses on AlpacaDatashort models. We divide the benchmarks into three groups based on the relationship between AlpacaDataShort model performance and PLW.
Group I benchmarks showed a negative quadratic relationship with the transformed PLW variable wp: ARC Challenge, PIQA, TruthfulQA-Gen, and WinoGrande. See figure 4 for results.
Group II benchmarks showed an increasing relationship between performance and wp on AlpacaDataShort: TruthfulQA-MC2, Alpaca Eval 1 (Mixtral), and PandaLM. Interestingly, Alpaca Eval 1 (Mixtral) and PandaLM are both long generation benchmarks, but the generation version of the TruthfulQA benchmark is in the first group while the multiple choice benchmark is in this group.
Group III are high noise benchmarks, limited to the six WMT translation tasks. See figure 6 for individual plots. In general WMT tasks showed very small variance, at most varying across only a few percentage points. To reduce the noise, these benchmarks were calculated over both the validation and test sets, but noise still precludes making clear conclusions about performance on any benchmark. Of note, combining benchmarks that translate from English produces a trend line suggestive of a positive relationship with wp. See figure 5(a) for an example. Figure 5(b) portrays into English translation benchmarks, but no trends could be identified.
(a) ARC Challenge
(b) PIQA
(c) TruthfulQA-Gen
(d) WinoGrande
Figure 4: Group I benchmark performance. Note the negative quadratic relationship with transformed PLW.
(a) TruthfulQA-MC2
(b) Alpaca Eval (AE) v1
(c) PandaLM
Figure 5: Group II benchmarks showed increasing performance with PLW.
(a) Translation From English
(b) Translation To English
[…]
Figure 6: Group III benchmarks showed little relationship between performance and PLW.
Appendix B Recommended Prompt Loss Weight
In this section, we present additional figures and tables related to optimal prompt loss weights (PLWs) as predicted by fitted generalized additive models (GAMs).
The optimal PLW predictions from the paper body were based on all benchmarks. Here we present three different relative aggregates: all benchmarks (labeled “All”), multiple choice benchmarks (labeled “MC”), and non-translation generation benchmarks (labeled “Gen”). We found that the translation benchmarks contributed little to the predictive power of the fitted GAMs. In an attempt to predict PLW values in a more relevant context, we did not include the translation benchmarks in the “Gen” model. However, they are still included in the “All” model to capture a wider range of tasks.
See figures 7,8,9 for contour plots for the “All”, “MC”, and “Gen” benchmarks, respectively. The “All” contour plot is identical to the contour plot in the main body.
See table 4 for a list of optimal PLWs over a range of completion-prompt ratios. Note that the predicted optimal PLWs confirm the conclusions from our regression analysis in section 5.2. Namely that PLW is important for short-completion data, but can be ignored for long-completion data.
$R_{g}$ Optimal PLW
All MC Gen
8.0 1.000* 1.000* 0.654
7.5 1.000* 1.000* 1.000*
7.0 1.000* 1.000* 1.000*
6.5 1.000* 1.000* 1.000*
6.0 1.000* 1.000* 0.000*
5.5 1.000* 1.000* 0.000*
5.0 0.000* 1.000* 0.000*
4.5 0.000* 1.000* 0.000*
4.0 1.000* 1.000* 0.000*
3.5 1.000* 1.000* 0.000*
3.0 1.000* 1.000* 1.000*
2.5 1.000* 1.000* 1.000*
2.0 0.239 1.000* 0.679*
1.5 0.183 0.278 0.385*
1.0 0.155 0.183 0.321*
0.5 0.155 0.155 0.292*
Table 4: Optimal prompt loss weight (PLW) per completion-prompt length ratio $R_{g}$ on all benchmarks (“All”); multiple choice benchmarks (“MC”); and the combination of TruthfulQA-Gen, Alpaca Eval 1, and PandaLM benchmarks (“Gen”). Predictions are based on the ratio-PLW interaction term of fitted generalized additive models.
*The difference between the maximum and minimum predicted values at this ratio is less than 5% of the score range.
Figure 7: “All” Ratio-PLW Interaction for all benchmarks.
Figure 8: “MC” Ratio-PLW Interaction for multiple choice benchmarks.
Figure 9: “Gen” Ratio-PLW Interaction for non-translation generation benchmarks.
Appendix C Reproducibility
This section provides technical details on all experiments and benchmarks for transparency and to encourage reproduction of results. To help with reproducibility, we will also upload the fine-tuned models, test generation outputs, and the AlpacaDataShort dataset to the HuggingFace Hub.
C.1 Model Fine-Tuning
Model fine-tuning was performed with the Stanford Alpaca GitHub repository at www.github.com/tatsu-lab/stanford_alpaca/tree/761dc5b.
To experiment with prompt loss weight, we modified HuggingFace’s Transformers library to allow specifying a loss_weights parameter for LlamaForCausalLM’s forward method.
We used the following commit of Transformers https://github.com/huggingface/transformers/tree/3b7675b.
All models were trained on a single four A100 80GB node and we used the first set of hyperparameters recommended in the Fine-tuning subsection of Stanford Alpaca’s README.md file, except for the three experimental variables: pretrained model, prompt loss weight, and training dataset.
AlpacaData is available from the Stanford Alpaca repository. AlpacaDataCleaned can be found at www.github.com/gururise/AlpacaDataCleaned/tree/791174f and is labeled “alpaca_data_cleaned.json”. As noted above, we will release AlpacaDataShort in the near future to the HuggingFace Hub.
C.2 Model Evaluation
We used three evaluation frameworks: EleutherAI’s Language Model Evaluation Harness (EEH), AlpacaEval 1, and PandaLM.
In an effort to match the current HuggingFace Open LLM leaderboard, we evaluated ARC Challenge, TruthfulQA-MC2, WinoGrande, and PIQA on the same EEH commit that the HuggingFace leaderboard uses: www.github.com/EleutherAI/lm-evaluation-harness/tree/b281b09 We also matched the number of shots with the number used for the HuggingFace leaderboard for ARC Challenge, TruthfulQA-MC2, and WinoGrande.
TruthfulQA-Gen and all translation tasks were evaluated using a more recent commit at www.github.com/EleutherAI/lm-evaluation-harness/tree/b93c3bc. We modified the translation tasks at this commit to include an appropriate prompt to support zero-shot translation.
Though version 2 of AlpacaEval has recently been released, we used version 1 from the following commit https://github.com/tatsu-lab/alpaca_eval/tree/495b606. To use Mixtral 8x7B as an auto-evaluator for AlpacaEval 1, we modified the Guanaco-33b evaluator’s config and prompt minimally to match Mixtral’s format. Models were evaluated on the default test set which can be found at https://huggingface.co/datasets/tatsu-lab/alpaca_eval/blob/main/alpaca_eval.json. We plan on submitting a pull request with these additions in the near future.
For PandaLM, we used the commit at https://github.com/WeOpenML/PandaLM/tree/eb758c4 and evaluated on version 1 of the default test set (found at “data/testset-inference-v1.json” in the PandaLM repository).
C.3 Regression
All statistical analysis and regression modeling was performed with R, version 4.3.0. We used the glmmTMB library, version 1.1.8, to perform generalized linear mixed modeling (GLMM) and validated results with the same library and with DHARMa, version 0.4.6.
C.4 Causal Mechanism
Most of the analysis performed to shed light on the causal mechanism should version and implementation agnostic. However, BLEU score implementations vary widely, and we used sacre BLEU to evaluate memorization of the training set. We used Corpus BLEU from the sacreBLEU library at https://github.com/mjpost/sacrebleu. Instead of a commit hash, we share the metric signature:
‘‘nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.4.0’’
C.5 Predictive Model
We fit several generalized additive models (GAMs) using the mgcv library, version 1.9-1 and the same version of R as above, version 4.3.0.
Appendix D Artifact Licensing
We respected all licenses for artifacts and resources used in this research. Please see table 5 for an overview of primary resources and licenses.
Resource License Application
Transformers Apache 2.0 Model Training
Stanford Alpaca Apache 2.0 Model Training
LLaMA 1 LLaMA License Pre-trained model weights
LLaMA 2 LLaMA 2 Community License Pre-trained model weights
Mixtral 8x7B Apache 2.0 Model Evaluation
EleutherAI’s LM Evaluation Harness MIT Model Evaluation
AlpacaEval 1 Apache 2.0 Model Evaluation
AlpacaEval Dataset CC BY-NC 4.0 Model Evaluation
PandaLM Apache 2.0 Model Evaluation
ARC Challenge CC BY-SA 4.0 Model Evaluation
PIQA AFL 3.0 Model Evaluation
TruthfulQA Apache 2.0 Model Evaluation
WinoGrande Apache 2.0 Model Evaluation
WMT 14 No License Model Evaluation
WMT 16 No License Model Evaluation
Table 5: Licenses for resources used in this research.

### v3

@@
 Abstract
-We present a study analyzing the effects of prompt loss weighting (PLW) on supervised instruction fine-tuning.
-We recreated Stanford’s Alpaca experiment with both LLaMA 1 and LLaMA 2 and multiple instruction datasets.
-We found that performance of models fine-tuned on our short-completion dataset had a statistically-significant negative quadratic relationship with PLW, but performance of models fine-tuned on medium- and long-completion data did not show any relationship with PLW.
-I.e., prompt loss can be safely ignored for many datasets.
-For short-completion data, small values ($0.01-0.1$) of PLW were optimal for multiple-choice and short-generation tasks while large values ($\approx 1.0$) of PLW were optimal for long-generation tasks.
-We concluded that low non-zero PLW encourages models to not diverge from pre-trained model weights during training and high PLW reduces overfitting.
-Finally, we present a rough guide for selecting PLW values based on the completion-prompt length ratio of fine-tuning data.
+We present a novel study analyzing the effects of various prompt loss token weights (PLW) for supervised instruction fine-tuning (SIFT).
+While prompt-masking (PLW $=0$) is common for SIFT, some fine-tuning APIs support fractional PLWs and suggest that using a small non-zero PLW can help stabilize learning when fine-tuning on short-completion data.
+However, there has never been a study confirming this claim, and OpenAI, a major cloud-based SIFT provider, recently removed this parameter from their fine-tuning API.
+We found that performance of models fine-tuned on short-completion data had a statistically-significant negative quadratic relationship with PLW.
+Using small values ($0.01-0.5$) of PLW produced better results on multiple-choice and short-generation benchmarks (outperforming models fine-tuned on long-completion data) while large values ($\approx 1.0$) of PLW produced better results on long-generation benchmarks.
+We explained this effect and verified its importance through additional experiments.
+This research serves as a warning to API providers about the importance of providing a PLW parameter for SIFT.
 Instruction Fine-Tuning: Does Prompt Loss Matter?
-Mathew Huerta-Enochian EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul mathew@eq4all.co.kr
+Mathew Huerta-Enochian  and Seung Yong Ko EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul {mathew, stephenko}@eq4all.co.kr
 1 Introduction
 Recent research in language modeling has made huge advances in training instruction-following agents.
-Both supervised fine-tuning and reinforcement learning have been employed to much success.
-However, our understanding of optimal hyperparameters and standards of practice have been slow to catch up.
-This research intends to contribute to hyperparameter selection for supervised instruction fine-tuning (SIFT) via an in-depth analysis of a single training hyperparameter: prompt loss weight (PLW).
-The goal of SIFT is to teach a model to perform a language task given a prompted instruction.
+Both supervised fine-tuning (SFT) and reinforcement learning (RL) have been employed to much success.
+However, our understanding of optimal hyperparameters and standards of practice (SOPs) have been slow to catch up.
+This research contributes to supervised instruction fine-tuning (SIFT) SOPs via an in-depth analysis of a single training hyperparameter: prompt loss weight (PLW).
 While training, model parameters are updated by optimizing for next-token maximal likelihood classification.
-However, it is unclear if this optimization should be performed on the entire prompt-completion sequence or only on the target completion.
-Most open sourced solutions for supervised fine-tuning (SFT) either mask the prompt loss or use the entire sequence loss.
-On the other hand, OpenAI previously supported a prompt_loss_weight parameter that allowed users to specify PLW during SIFT.
-We make the following contributions related to this parameter:
-  • We show that PLW has a statistically significant negative quadratic relationship with model performance when fine-tuning on short-completion instruction data.
-  • We show that PLW and prompt masking parameters can be safely ignored when fine-tuning on medium- or long-completion data, streamlining the fine-tuning process.
-  • We provide evidence that the benefits of PLW for short-completion data are due to a regularizing effect and not due to previously suspected mechanisms.
-  • We present a baseline for PLW selection for short-completion fine-tuning.
-We provide relevant background and hypotheses in section 2 and section 3, respectively.
-We resent our methodology in section 4.
-Descriptive and regression analysis and an exploration of causal mechanisms are presented in section 5.
-We extend the experiment to additional datasets and present a basic model for prompt loss weight selection in section 6.
-We present conclusions in section 7.
+Most open sourced solutions for SFT either mask the prompt loss (for prefix language modeling) or use the entire sequence loss (for full language modeling) while some API providers support an explicit PLW parameter that allows users to apply fractional PLW during SIFT.
+The commonly-held notion is that fractional PLW helps stabilize when fine-tuning on data with short outputs.
+Recently, however, OpenAI quietly removed support for their prompt_loss_weight parameter.
+The reason for the removal is unknown. Furthermore, to our knowledge, there has never been a proper study on the effects of PLW.
+We make the following contributions:
+  • We showed that PLW has a significant relationship with model performance when fine-tuning on short-completion data.
+  • We showed that this relationship is due to a combination of regularizing effects and not the accepted explanation of increased training stability.
+  • We provided evidence that PLW cannot be replaced by other common regularizers alone and that PLW is important for strong performance on short-generation downstream tasks.
+  • We verified that PLW can be safely ignored when fine-tuning on long-completion data.
+We provide relevant background and hypotheses in section 2 and 3, respectively.
+The main regression experiment is presented in sections 4 and 5.
+Supplemental experiments further validating our claims are presented in section 6.
+We present conclusions in section 7 followed by several appendices for additional analysis and discussion.
 2 Background
 2.1 Definitions
-We define instruction data as one or many instances of structured text data, each containing an instruction text, an optional context or input text, and a target output text.
-For the rest of the paper, we use the term prompt to refer to the concatenation of instruction and input texts and the term completion to refer to the target output text.
-SIFT attempts to train a model to generate an appropriate completion for a given prompt.
+We define instruction data as one or many instances of structured text data, each containing an instruction, an optional input, and a target output text.
+We will use the term prompt to refer to the concatenation of the instruction and input (if it exists) and the term completion to refer to the target output.
+The goal of SIFT is to fine-tune a model to generate an appropriate completion for a given prompt.
 We define the generation ratio $R_{g}$ as the ratio of completion length to prompt length.
@@
 HuggingFace’s Transformers library Wolf et al.
-(2020), the de facto library for training LLMs, allows users to mask select tokens when calculating loss.
-In Transformers, optimizing next-token prediction on prompt tokens is therefore binary—either token loss is masked (PLW $=0$) or it is unmasked (PLW $=1$).
-Until recently, OpenAI supported a prompt_loss_weight parameter in their fine-tuning API, but it was officially removed as part of the v1 fine_tune API deprecation in early January, 2024.
-This prompt_loss_weight parameter used a default value of $0.01$ with the following parameter explanation:
-“This controls how much the model tries to learn to generate the prompt (as compared to the completion which always has a weight of 1.0), and can add a stabilizing effect to training when completions are short.
+(2020), the de facto library for training LLMs, allows users to mask select tokens when calculating token classification loss.
+In Transformers, weights for next-token prediction loss is therefore binary—either token loss is masked (PLW $=0$) or it is unmasked (PLW $=1$).
+As mentioned in section 1, OpenAI officially removed support for a prompt_loss_weight parameter in their fine-tuning API as part of the v1 fine_tune API deprecation in early January, 2024.
+This prompt_loss_weight parameter used a default value of $0.01$ with the following parameter explanation: “This controls how much the model tries to learn to generate the prompt (as compared to the completion which always has a weight of 1.0), and can add a stabilizing effect to training when completions are short.
 If prompts are extremely long (relative to completions), it may make sense to reduce this weight so as to avoid over-prioritizing learning the prompt.”
-Though we could not find a study validating OpenAI’s claim or any literature that presents an analysis of PLW, there were several studies that reported parameter values for prompt_loss_weight when fine-tuning OpenAI models.
+Though we could not find a study validating OpenAI’s claim or any literature that presents an analysis of PLW, we found several studies that reported using this parameter.
 Though they do not provide their reasoning, Kozachek (2023) reported that they fine-tuned GPT-3 with a prompt_loss_weight of $0.1$.
@@
 AlpacaDataCleaned 14.21 (9.95) 6.42 (17.65) 162.74 (150.89) 9,490,837 7.83
-AlpacaDataShort 16.93 (13.10) 162.34 (151.69) 14.62 (10.99) 10,035,667 0.082
+AlpacaDataShort 16.93 (13.10) 162.34 (151.69) 14.62 (10.99) 10,035,667 0.08
+UltraFeedback* 184.47 (268.04) 0 (-) 327.37 (291.48) 31,075,393 1.77
+DatabricksDolly* 18.65 (74.55) 91.60 (250.66) 91.14 (149.15) 3,023,113 0.83
+UltraFeedbackShort* 94.09 (141.94) 206.18 (211.03) 55.03 (61.00) 16,351,333 0.18
+DatabricksDollyShort* 18.83 (75.29) 160.37 (270.17) 28.79 (47.49) 3,122,209 0.16
 Table 1: Dataset statistics: mean tokenized instruction, input, and completion sequence lengths (standard deviations in parentheses), total token counts for each dataset, and the generation ratio $R_{g}$.
+* Used for supplemental experiments in section 6.
 3 Hypotheses
-Based on OpenAI’s documentation for prompt_loss_weight, we expect PLW to have a stabilizing effect on LLM SIFT when using short-completion data.
-Our expectation was that PLW preserves the hidden representation learned for next-token-classification when most embeddings are being used to predict tokens at the end of the sequence (as is the case when fine-tuning on short-completion data).
-This would represent a positive relationship between PLW and downstream performance.
-However, unless the model is expected to generate prompt-like data in downstream tasks, the training and downstream sequence contexts will not match.
-In other words, training a model to maximize next-token-prediction on prompt tokens should be most useful for generating instruction data.
-Due to this train-test mismatch, we expect PLW to also have a negative influence on downstream performance.
-Based on the assumption that some amount of PLW is beneficial for short-completion SIFT, we expect these two competing factors to result in a downward curved relationship between downstream performance and PLW.
+Based on OpenAI’s explanation for prompt_loss_weight, we expected that for SIFT with short-completion data and small values of PLW, there would be a positive relationship between PLW and downstream performance.
+However, training a model to maximize next-token-prediction on prompt tokens should be most useful for generating instruction data, and over-prioritizing prompt token loss should have a negative influence on downstream performance.
+Based on these assumptions, we would expect the two competing factors to result in a downward curved relationship between PLW and downstream performance.
 Limiting PLW to the range of [0, 1], we postulate that there is a critical value $\lambda$ for PLW with $10<=\lambda<=$ 1.
 For PLW less than $\lambda$, the positive effect dominates the negative effect and for values greater than $\lambda$, the negative effect dominates the positive effect.
-If $\lambda=0$, then PLW’s contribution to model performance is strictly negative.
-Conversely, if $\lambda=1$, then PLW contributes strictly positively to model performance.
-We postulate that both $\lambda$ and the magnitude of the effect are inversely proportional to $R_{g}$ since there will be $R_{g}$ as many completion token loss terms than prompt loss terms.
+If $\lambda=0$, then PLW’s contribution to model performance is strictly negative, and if $\lambda=1$, then PLW contributes strictly positively to model performance.
 Note that $\lambda$ is not an intrinsic characteristic of the dataset, model architecture, or training algorithm.
 Rather, it would depend on numerous factors and change for each task.
-We tested the relationship between PLW and model performance using three fine-tuning datasets spanning a range of $R_{g}$ values: 0.327, 7.83, and 0.082.
-We present the following hypotheses to be tested against each dataset.
+We then made two hypotheses to test the above relationship and performed regression analysis on three fine-tuning datasets spanning a range of $R_{g}$ values: 0.08, 3.27, and 7.83.
 Null Hypothesis ($\mathbf{H_{0}}$) Prompt loss weight has no relationship with model performance.
 Alternative ($\mathbf{H_{1}}$) Prompt loss weight has a quadratic relationship with model performance.
-We fixed the threshold $\alpha=0.05$ significance level, and we expected to be able to reject $\mathbf{H_{0}}$ for models trained on short-completion data.
+We used the standard $\alpha=0.05$ significance level, and we expected to reject $\mathbf{H_{0}}$ only for models trained on short-completion data.
 4 Methodology
@@
 Training was performed exactly as per the original Alpaca experiment, and we used the hyperparameters suggested by the authors, modifying only the three experimental parameters (PLW, PTLM, dataset) with each run.
-We provide additional details for reproducibility in appendix C and will release our trained models on the HuggingFace Hub.
+We provide additional details for reproducibility in appendix E and will release our trained models on HuggingFace’s Hub.
 4.1 Prompt Loss Weight
@@
 4.3 Fine-Tuning Dataset
-We ran all experiments with three datasets: AlpacaData (the instruction dataset from the original Alpaca experiment), AlpacaDataCleaned, and AlpacaDataShort.
-AlpacaDataCleaned Ruebsamen (2023) is a cleaned and curated version of AlpacaData that has recently been combined with data from the GPT4 LLM dataset Peng et al. (2023).
+We ran all experiments with three datasets: AlpacaData (the instruction dataset from the original Alpaca experiment), AlpacaDataCleaned Ruebsamen (2023), and AlpacaDataShort.
+AlpacaDataCleaned is a cleaned and curated version of AlpacaData that has recently been combined with data from the GPT4 LLM dataset Peng et al.
+(2023).
 Cleaning is noted as ongoing and includes fixes for the following issues in AlpacaData: hallucinations, merged instructions, empty outputs, empty code examples, instructions to generate images, N/A outputs, inconsistent input fields, wrong answers, nonsensical instructions, and extraneous control characters.
-We generated AlpacaDataShort from AlpacaDataCleaned by swapping instructions and completions for all long-completion instances and rephrasing the task as an instruction-prediction task.
-For instances with empty input, we used
-Predict the prompt that generated the following AI output.
-as the new instruction field. For instances with non-empty input, we used
-Predict the prompt that generated the below AI output given the following context. Context: <original_input>
-as the new instruction field. In all cases we then used the original completion as the new input field and the original instruction as the new completion field.
-Descriptive statistics for tokenized instruction, input, and completion sequences are presented in table 1.
+We generated AlpacaDataShort from AlpacaDataCleaned by rephrasing long-completion instances as prompt-prediction task, a process we denote as prompt inversion.
+See Appendix A for more on prompt inversion.
+Descriptive statistics for these each dataset are presented in table 1.
 Note that AlpacaDataCleaned is strongly long-completion with an $R_{g}$ of $7.83$ while AlpacaDataShort is short-completion with an $R_{g}$ of $0.082$.
@@
 PandaLM 1 0 Test GL
-Table 2: Benchmark tasks used for evaluation.
-All benchmarks were scored using Eleuther AI’s Language Model Evaluation Harness, the AlpacaEval 1 framework, and the PandaLM framework.
+Table 2: Evaluation benchmarks.
 Validation splits were used when test splits were unavailable, and validation and test splits were combined for noisy benchmarks.
@@
 †This benchmark was calculated with num_fewshot $=0$ but uses a built-in minimum of 6 shots.
-We evaluated each model on thirteen instruction benchmarks covering both multiple choice and text generation tasks.
-We attempted to select benchmarks that were relatively cheap to compute and covered a range of tasks.
+We evaluated each model on thirteen instruction benchmarks covering multiple choice and text generation tasks.
+We selected benchmarks that were relatively cheap to compute and covered a range of tasks.
 We used three evaluation frameworks: EleutherAI’s Language Model Evaluation Harness (EEH) Gao et al.
-(2023), Alpaca Eval 1 Li et al.
+(2023b), AlpacaEval 1 Li et al.
 (2023), and PandaLM Wang et al.
@@
 Eleven benchmarks were run using EEH.
-Four benchmarks weremultiple choice tasks: ARC Challenge Clark et al.
+Four of these were multiple choice tasks: ARC Challenge Clark et al.
 (2018), PIQA Bisk et al.
@@
 (2019).
-Seven benchmarks were short generation tasks: TruthfulQA-Gen Lin et al.
+Seven were short generation tasks: TruthfulQA-Gen Lin et al.
 (2022) and six WMT14 and WMT16 translation benchmarks Bojar et al.
-(2014, 2016), limited to four languages the PTLMs learned during pretraining (English, French, German, and Romanian).
-For WMT benchmarks, we used the instruction “Translate the following from <src_lang> to <tgt_lang>” and evaluated over both the validation and test sets to reduce variance.
+(2014, 2016), limited to four languages the PTLMs saw during pretraining (English, French, German, and Romanian).
+For WMT benchmarks, we used the zero-shot instruction “Translate the following from <src_lang> to <tgt_lang>” and evaluated over both the validation and test sets to reduce variance.
 Long-generation performance was evaluated using AlpacaEval 1 and PandaLM, which are both LLM-as-a-judge frameworks.
-The default auto-evaluator for Alpaca Eval 1 is GPT4, but evaluation of all training runs would be beyond the scope of this research, so we used Mixtral 8X7B Jiang et al.
+The default auto-evaluator for AlpacaEval 1 is GPT4, but using paid APIs would beyond the scope of this research, so we used Mixtral 8X7B Jiang et al.
 (2024) as an auto-evaluator.
-Mixtral performed the best of all available LLMs that we tested on Alpaca Eval’s evaluator test dataset Dubois et al.
-(2023), scoring 64.9 for agreement with human evaluators, just below Claude’s Anthropic (2023) 65.3 agreement.
-For reference, the default GPT4 auto-evaluator scored a 70.99 human agreement.
+Mixtral performed the best of all open-source LLMs that we tested on AlpacaEval’s evaluator test dataset Dubois et al.
+(2023), with 64.9% agreement with human evaluators.
+For reference, Claude Anthropic (2023) has 65.3% and GPT4 has 70.99% human agreement.
 5 Results and Discussion
@@
 Best viewed in color.
+A visualization of the simple and min-max scaled relative performance aggregates by wp and dataset is presented in figure 1.
+Note that all analysis was performed using the relative aggregate.
+The simple aggregate is dominated by large performance changes on a few benchmarks and is included only for completeness.
 5.1 Performance Trends
-A visualization of the simple and relative performance aggregates by wp and dataset is presented in figure 1.
-The simple aggregate is the unweighted mean of all task scores for given wp and dataset values.
-For the relative aggregate, we first min-max normalized benchmark scores over each dataset-benchmark group and then performed aggregation.
-The simple aggregate is dominated by large performance changes on the AlpacaEval 1 and PandaLM benchmarks.
-Therefore, all analysis was performed using the relative aggregate and the simple aggregate was visualized only for completeness.
-[…]
-Identification of trends on the six translation benchmarks was less clear due to the high level of noise relative to performance range.
-Aggregating scores from benchmarks translating from English to the paired language benchmarks is suggestive of an increasing relationship, but is incredibly noisy.
-The performance difference across wp values for the two long-generation benchmarks was around twenty percentage points, in stark contrast to the less than two percentage point change for short-generation and multiple choice benchmarks.
-This suggests that PLW plays an important role in the ability to generate high quality text, and the optimal PLW for short-generation and long-generation benchmarks is clearly different.
-Finally, performance of LLaMA 2 models was higher than that of LLaMA 1 models and performance of AlpacaDataCleaned models were higher than that of AlpacaData models, validating the improvements of LLaMA 2 and AlpacaDataCleaned over their predecessors.
-See Appendix A for detailed figures for each benchmark.
+There are several qualitative performance trends that are of interest. For more thorough discussion and task-specific benchmark plots, see Appendix B.
+A group of four benchmarks (Arc Challenge, PIQA, TruthfulQA-Gen, and WinoGrande) clearly show the expected negative quadratic relationship between wp and performance of AlpacaDataShort models.
+Notably, AlpacaDataShort models outperform AlpacaData and AlpacaDataCleaned on this group given optimal PLW tuning.
+On the long-generation benchmarks, however, AlpacaDataShort models show a steadily-increasing trend.
+For these seven benchmarks and AlpacaDataShort models, using prompt loss masking (wp$=0$) almost always produced the worst scores.
+The wp-based performance increase was around twenty percentage points for both long-generation benchmarks.
+Contrasted with the less than two percentage point change for short-generation and multiple choice benchmarks, and it is evident that the optimal PLW for short-generation and long-generation benchmarks is different.
+Clear qualitative performance trends for the AlpacaData and AlpacaDataCleaned and for any model on the six translation benchmarks could not be identified.
 5.2 Regression
-For each data group, we fit a generalized linear mixed model (GLMM) predicting min-max transformed benchmark scores.
-As with the min-max transformation for the relative aggregate, benchmark scores were min-max transformed over dataset-benchmark groups.
+For each data group, we fit a generalized linear mixed model (GLMM) with the relative aggregate benchmark scores as the response variable.
 We expected a quadratic relationship between the score and wp, so we included a second order polynomial of wp as a fixed effect.
-Furthermore, we knew that the performance-PLW relationship varies by benchmark and that scores were min-max normalized over each benchmark, so we used a random slope with respect to benchmark.
+Furthermore, we knew that the PLW-performance relationship varies by benchmark and since scores were min-max normalized over each benchmark, we used a random slope (and no intercept) with respect to benchmark.
 Since we did not min-max normalize over PTLM groups and since we saw consistent improvement when using LLaMA 2, we modeled a random intercept for PTLM.
@@
 For the AlpacaDataShort case, min-max transformed performance showed a statistically significant negative quadratic relationship with wp at our target $\alpha=0.05$ significance level.
-This means that while wee could not reject the null hypothesis for the AlpacaData and AlpacaDataCleaned scenarios, for the AlpacaDataShort scenario, there was sufficient evidence to reject the null hypothesis in favor of the alternative hypothesis H1.
+This means that while we could not reject the null hypothesis for the AlpacaData and AlpacaDataCleaned scenarios, for the AlpacaDataShort scenario, there was sufficient evidence to reject the null hypothesis in favor of the alternative hypothesis H1.
 Using the fixed effect coefficients, we can predict the critical PLW value $\lambda$ for AlpacaDataShort fine-tuning that maximizes the min-max transformed benchmark scores.
@@
 5.3 Causal Mechanism & Interpretation
-To identify possible causal mechanisms, we first investigated the effects of PLW on training stability by analyzing training loss relative standard deviation (RSD) over all five-step windows.
+To identify possible causal mechanisms, we first investigated the effects of PLW on training stability by analyzing training loss relative standard deviation (RSD) over five-step windows.
 See figure 2(a) for a boxplot of mean RSD for each model.
 For all dataset and PTLM factors, increasing PLW from zero led to a sharp increase in mean RSD and then a slow decrease to a minimum mean RSD at PLW $=1$.
-There is no obvious explanation for why training loss RSD would initially increase before decreasing.
-If training loss stability was the primary factor in improved performance, we would expect stability to be best for PLW between $0.01$ and $0.1$ and for performance at PLW $=0$ to be similar with performance at PLW $=0.01$ since mean RSD at these values are similar.
+There is no obvious explanation for why training loss RSD would increase for small PLW before decreasing for large PLW.
+If training loss stability was the primary factor in improved performance, we would expect RSD to be lowest for PLW between $0.01$ and $0.5$ (or even between $0.01$ and $0.1$ based on short-generation benchmarks) and for performance at PLW $=0$ to be similar with performance at PLW $=0.01$ since mean RSD at these values are similar.
 However, mean RSD drops by a factor of two across the PLW $\in[0.01,0.1]$ range, and performance at PLW $=0.01$ is significantly higher than the masked prompt loss scenario.
@@
 We would expect weights to change more when loss is erratic, but the range of PLW values that better preserved PTLM weights was similar to the range of increased training loss RSD.
-Unfortunately, we do not have an explanation for this apparent contradiction and simply conclude that PTLM model weights were better preserved for small non-zero PLW despite high loss instability.
+This is an interesting result, and we conclude that PTLM model weights were better preserved for small non-zero PLW despite high loss instability.
 We next explored how PLW affected training data memorization.
@@
 Corpus BLEU then decreased as PLW increased from 0.1.
-We also analyzed generation length on the AE1 test set which showed high deviations in generation length for the high loss RSD PLW values for LLaMA 2 only and a generally increasing length trend with PLW.
+We also analyzed generation length on the AlpacaEval 1 test set which showed high deviations in generation length for the high loss RSD PLW values for LLaMA 2 only and a generally increasing length trend with PLW.
 Considering that AlpacaDataShort contains short completions, this strongly suggests that PLW contributes a regularizing effect and helps decrease overfitting.
-The positive relationship between performance and PLW on Alpaca Eval and PandaLM as opposed to the negative quadratic relationship seen on several of the other benchmarks can then be explained by a tradeoff between longer generations and decreased overfitting for high PLW and better preservation of PTLM weights for low PLW.
-6 Predicting Prompt Loss Weight
-Figure 3: Ratio-PLW Relationship
-We have shown that PLW is an important hyperparameter when fine-tuning on short-completion data but is effectively irrelevant when using long-completion data.
-We now present a model for predicting an optimal PLW given a dataset’s $R_{g}$.
-We first extended the previous experiments to two additional datasets: AlpacaDataMedium and AlpacaDataTiny to increase coverage of the parameter space.
-AlpacaDataMedium and AlpacaDataTiny have $R_{g}$ values of 1.0 and 0.042, respectively, and were generated similarly with AlpacaDataShort, but with different target completion-prompt ratios.
-We then fit several generalized additive model (GAMs) with a tensor smooth for the interaction between PLW and $R_{g}$.
-GAMs offer more flexible modeling than linear models but at the expense of interpretability.
-We fit our models using the R library mgcv and the following equation:
-score ~ te(w, r, k=3) + factor(b)”, where te is a full tensor product smooth, w is the untransformed PLW parameter, r is the $R_{g}$, k is the number of splines, and b is the benchmark task.
-Using the fitted w-r interaction, we can now estimate an optimal PLW value for a given completion-prompt ratio $R_{g}$. See figure 3 for a visualization.
-Roughly, the fitted interaction term recommends using PLW $=0.155$ for small completion-prompt length ratios ($R_{g}\leq 1$) and up to PLW $=0.278$ for a $R_{g}=1.5$ for optimal performance across all tasks.
-This prediction is close to our regression-predicted value of 0.242.
-The interaction term also confirms our conclusion that PLW is less important for data with relatively long completions.
-This is a rough guide, and we recommend using it as a starting point. See appendix B for additional plots using different benchmark tasks.
+The positive relationship between performance and PLW on AlpacaEval 1 and PandaLM as opposed to the negative quadratic relationship seen on several of the other benchmarks can then be explained by a tradeoff between longer generations and decreased overfitting for high PLW and better preservation of PTLM weights for low PLW.
+6 Supplemental Experiments
+In this section, we present supplemental PLW experiments that suggest that PLW cannot be replaced by alternative regularization techniques for SIFT and that the effects of PLW extend to other short-completion datasets.
+Since the translation benchmarks showed high levels of noise and unclear correlation with PLW in the main experiment, they were not included in supplemental experiments.
+6.1 Alternative Regularizers
+LLaMA Regularization Aggregate
+Simple Relative
+1 Weight Decay 0.507 0.795
+Minkowski Metric 0.505 0.812
+Dropout 0.500 0.783
+Label Smoothing 0.500 0.785
+PLW (Ours) 0.506 0.855
+2 Weight Decay 0.537 0.773
+Minkowski Metric 0.538 0.772
+Dropout 0.541 0.805
+Label Smoothing 0.538 0.837
+PLW (Ours) 0.537 0.894
+Table 4: Supplemental comparison of AlpacaDataShort models fine-tuned with various regularizers.
+High scores are in bold.
+As explained in section 5, the relative aggregate should be used for analysis, and the simple aggregate is provided for reference only.
+To investigate if the effects of PLW on short-completion SIFT can be emulated with other common regularization techniques, we repeated the AlpacaDataShort training runs for PLW $=0$ and PLW $=1$ while applying various regularizers.
+We tested weight decay, minimizing the Minkowski distance between PTLM weights and learned weights, dropout, and label smoothing.
+See Appendix C.1 for additional visualizations.
+Note that we do not compare PLW with KL-divergence.
+While using KL-divergence as a regularizing loss is common for SFT in general and is used in RL-based LLM alignment (Stiennon et al., 2020; Korbak et al., 2022; Gao et al., 2023a), collecting embeddings for any non-trivial dataset to be used for LLM SIFT presents a huge computation and memory overhead.
+As can be seen in the results in table 4, relative aggregate scores were higher for models fine-tuned with fractional PLWs.
+This suggests that PLW provides a unique regularizing effect that cannot be easily replaced with other regularizers.
+6.2 Alternative Datasets
+(a) Original Datasets
+(b) Prompt-Inverted Datasets
+Figure 3: Relative aggregate scores showing the effects of PLW for SIFT on alternative datasets.
+(a) UltraFeedbackCleaned and DatabricksDolly models.
+(b) UltraFeedbackShort and DatabricksDollyShort models.
+We repeated our main training experiment using two additional datasets, a cleaned and binarized version of UltraFeedback (Cui et al.
+(2023); denoted UltraFeedbackCleaned) and databricks-dolly-15k (Conover et al., 2023; denoted DatabricksDolly).
+We also trained on prompt-inverted versions of both datasets, denoted UltraFeedbackShort and DatabricksDollyShort, respectively, and expanded analysis to include LLaMA 3 8B AI@Meta (2024) as a PTLM.
+See length statistics for these four datasets in table 1 and appendix C.2 for additional visualizations.
+The combined relative aggregate scores for UltraFeedbackCleaned and DatabricksDolly and for UltraFeedbackShort and DatabricksDollyShort are visualized in figure 3.
+Visual inspection suggests that PLW affected learning for both groups of datasets.
+For the shortened variants, performance appears to have the same negative quadratic relationship with PLW as AlpacaDataShort models do in the main experiment.
+The relationship between PLW and performance for models fine-tuned on the original datasets is less clear but appears to have a generally increasing trend, unlike the long-completion datasets in the main experiment.
+We believe this is due to both UltraFeedbackCleaned and DatabricksDolly having much smaller completion-prompt length ratios $R_{g}$ than the AlpacaData and AlpacaDataCleaned datasets.
 7 Conclusion
-In this study, we explored the downstream effects of prompt loss weight (PLW) for LLM supervised instruction fine-tuning (SIFT).
-We showed that performance is unaffected for the two long-completion datasets (AlpacaData, the original Alpaca dataset, and AlpacaDataCleaned, an updated version of AlpacaData) while performance had a negative quadratic relationship with prompt loss weight on our short-completion dataset (with a completion-prompt ratio $R_{g}\approx 0.08$).
-Interestingly, the highest scores on the ARC Challenge, PIQA, and WinoGrande benchmarks were all from models trained on short-completion dataset.
-This suggests that given a good choice of prompt loss weight, long-completion training data may not be necessary for instruction fine-tuning LLMs for language understanding and short generation tasks.
-Conversely, using long-completion data should be more robust as prompt loss weight and masking parameters can be ignored.
-We further analyzed the AlpacaDataShort models and found that low non-zero PLW discourages deviation from the PTLM weights and high PLW reduces overfitting and encourages longer generation.
-We suggest that these two factors are responsible for the negative quadratic relationship between performance and PLW.
-Finally, we estimated optimal PLWs over a range of completion-prompt length ratios.
+In this study, we explored the effects of prompt loss weight (PLW) on LLM supervised instruction fine-tuning (SIFT).
+We found that PLW has a statistically significant effect on learning for our short-completion dataset, and proper tuning of PLW allowed short-completion-trained models to outperform long-completion-trained models on short-generation benchmarks.
+We showed that the causal mechanism was due to a balance between two different regularizing effects and not due to increased training stability as is commonly attributed.
+We showed that the measured relationship extends to additional SIFT datasets and that the effects could not be sufficiently emulated with other regularizers.
+Based on the above conclusions, we assert the following two points.
+  1. 1.
+    Since models fine-tuned on short-completion datasets and with properly tuned PLWs outperformed all other models on short-generation benchmark tasks, we conclude that PLW is critical for effectively fine-tuning for downstream short-generation tasks.
+  2. 2.
+    Given the importance of PLW and given that many SIFT datasets and almost all natural language understanding (NLU) datasets are short-completion datasets, we warn SFT API providers about the need for a PLW parameter to adequately cover a full range of modeling applications.
 Limitations
@@
   2. 2.
-    We used a fixed seed $=42$ for all experiments.
-Since we used pre-trained models and no layers were freshly initialized, there was little variance in initial experiments.
-Therefore, we decided to limit runs to a single seed.
+    Since we used pre-trained models and no layers were freshly initialized, there was little variance in initial experiments. We therefore limited runs to a single seed of $42$.
   3. 3.
-    Suggested values for PLW from section 6 are based on the included experiments.
+    Suggested values for PLW from section D are based on the included experiments.
 Best PLW values when fine-tuning different models or using different datasets or training regimes may vary from the relationships shown here, though we are still confident that performance will not vary significantly by PLW for long-completion data.
   4. 4.
-    LLM-as-evaluator approaches like PandaLM and Alpaca Eval 1 are still relatively new, and these approaches are still being actively developed.
-We chose to use Mixtral as an auto-evaluator for Alpaca Eval.
-Although this is not the recommended auto-evaluator, we felt its performance on the human-labeled test set was good enough to justify its use.
+    The focus of our research was on how PLW affected fine-tuning based on the completion-prompt ratio of the training dataset.
+However, the absolute length and size of the dataset will likely play a role in learning dynamics.
+It would be good to include that perspective in future research on token loss weights.
   5. 5.
-    A more confident conclusion about the causative mechanism could not be made.
-As with most techniques in machine learning, the suggested causative relationship should be validated through numerous experiments and projects.
-Unfortunately, since we expect most instruction fine-tuning datasets to be long-completion datasets, further research into PLW will likely be limited.
+    LLM-as-evaluator approaches like PandaLM and AlpacaEval 1 are still relatively new, and these approaches are being actively developed.
+We chose to use Mixtral 8x7B as an auto-evaluator for AlpacaEval 1 due to budget limitations.
+While we cite high human evaluation correlation with Mixtral and justify this decision in section 4.4, using the default auto-evaluator would be beneficial for better comparison with other research.
+  6. 6.
+    While we define short- and long-completion data as have a completion-prompt ratio $R_{g}$ lower and greater than $1$, respectively, we do not provide justification for using $1$ as the threshold.
+Choosing a meaningful reference would be helpful to future research.
 Ethical Considerations
@@
 We did not rely on human evaluators, and at no point in our research did we expose anyone to risk of harm.
-We do acknowledge that standard deep learning training methods have a high carbon footprint, and we performed a total of 100 fine-tuning training runs.
+We acknowledge that standard deep learning training methods have a high carbon footprint, and we performed over 200 fine-tuning training runs.
 Model outputs cannot be predicted in advance, and, while we release our model weights in the spirit of transparency and collaboration, models may hallucinate or produce offensive output.
-Our synthetic dataset was generated from another publicly released dataset and we did not perform additional content filtering.
-A warning about both of these issues will be released along with the model and dataset.
+Additionally, our shortened datasets were generated from publicly released data, and we did not perform additional content filtering.
+A warning about both of these issues will be released along with the models and datasets.
 Acknowledgements
-This research was supported by Gwangju City’s Artificial Intelligence Industry Cluster Agency (AICA) and by the Korean Ministry of Trade, Energy, and Industry (MOTIE) under the KoSIGN project [Grant Number: 20014406].
-This research was possible thanks to the support of my coworkers at EQ4ALL.
-Appendix A Benchmark Performance Figures
-This appendix contains individual benchmark performance plots and additional discussion of performance trends and results.
-Overall, models trained on AlpacaDataCleaned performed best on the selected benchmarks, with models trained on AlpacaDataShort meeting or outperforming these models on numerous benchmarks for optimal selection of PLW.
-The general PTLM trend was for LLaMA 2 models to outperform LLaMA 1 models, again with some exceptions.
-In general, models trained on AlpacaData and AlpacaDataCleaned did not show consistent relationships with PLW on the tested benchmark.
-The rest of this section focuses on AlpacaDatashort models. We divide the benchmarks into three groups based on the relationship between AlpacaDataShort model performance and PLW.
-Group I benchmarks showed a negative quadratic relationship with the transformed PLW variable wp: ARC Challenge, PIQA, TruthfulQA-Gen, and WinoGrande. See figure 4 for results.
-Group II benchmarks showed an increasing relationship between performance and wp on AlpacaDataShort: TruthfulQA-MC2, Alpaca Eval 1 (Mixtral), and PandaLM.
-Interestingly, Alpaca Eval 1 (Mixtral) and PandaLM are both long generation benchmarks, but the generation version of the TruthfulQA benchmark is in the first group while the multiple choice benchmark is in this group.
-Group III are high noise benchmarks, limited to the six WMT translation tasks.
-See figure 6 for individual plots.
-In general WMT tasks showed very small variance, at most varying across only a few percentage points.
-To reduce the noise, these benchmarks were calculated over both the validation and test sets, but noise still precludes making clear conclusions about performance on any benchmark.
-Of note, combining benchmarks that translate from English produces a trend line suggestive of a positive relationship with wp.
-See figure 5(a) for an example.
-Figure 5(b) portrays into English translation benchmarks, but no trends could be identified.
+This work was supported by the Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City (No.
+BA00000797, LLM-based sign language translation for weather forecasts).
+This research was possible thanks to the support of our coworkers at EQ4ALL.
+Appendices
+Appendix A Prompt Inversion
+Figure 4: Examples of modifying prompt-completion ratios using prompt inversion, best viewed in color.
+To prompt-invert instances, we re-frame the prompt-completion task as an original-prompt-prediction task.
+I.e., we teach the model to predict the original instruction given an example completion and optional input.
+In the first example above, prompt inversion changes the instance’s word-based completion-prompt ratio $R_{g}$ from $34/(7+0)=4.857$ to $7/(9+34)=0.163$.
+In order to experiment with short-completion fine-tuning datasets while also preserving a similar dataset to use as a reference, we propose using prompt inversion to rotate the instruction, input, and output fields of instances.
+Prompt inversion modifies an instance to use the original instruction as the new completion text and the original output as the new input text.
+The model is then given the following instruction:
+Predict the prompt that generated the following AI output.
+if the input field is empty and
+Predict the prompt that generated the below AI output given the following context. Context: <original-input>
+if there is an input field. See figure 4 for a visualization of this process.
+To generate short versions of our test instruction datasets, we use prompt inversion to modify every instance with a completion-prompt length ratio $R_{g}>1$, based on the tokenized lengths of each field and using prompt to mean the concatenation of the instruction and input fields as explained in section 2.1.
+Synthesis of the original input and output texts to predict the original instruction requires both language understanding and reasoning.
+It also does not train the model for long-generation (since new completions will generally be short) and thus does not prepare the model for train-of-thought generation that long-generation tasks intrinsically learn.
+Therefore, we propose that prompt-inverted instances are best suited for learning natural language understanding (NLU)-related tasks rather than general generation tasks.
+Appendix B Main Experiment Benchmarks
+This section presents additional qualitative analysis of benchmark performance and score visualizations for each benchmark.
+For both the simple aggregate and the relative aggregate, models trained on AlpacaDataShort showed a visual relationship with wp.
+Based on this visual relationship, we divide benchmarks into three groups.
+The first group showed a negative quadratic relationship with wp, with performance exceeding that of AlpacaDataCleaned models.
+This group consists of ARC Challenge, PIQA, TruthfulQA-Gen, and WinoGrande benchmarks, and optimal PLW values for these four benchmarks vary from PLW $=0.01$ to PLW $=0.1$.
+See figure 5 for individual benchmark visualizations.
+The second group of benchmarks showed steadily increasing performance as wp increased, before leveling off to maximum values near wp$=1$.
+This group is TruthfulQA-MC2, AlpacaEval 1, and PandaLM.
+It is surprising that TruthfulQA-MC2 shows a relationshp more similar to the long-generation benchmarks and TruthfulQA-Gen resembles the other multi-choice benchmarks.
+See figure 6 for individual benchmark visualizations.
+[…]
+The third group consists of the six translation benchmarks and showed unclear correlation between performance and wp.
+Though aggregating benchmarks into “to English” and “from English” subgroups creates visualizations suggestive of a relationship, benchmarks from this group showed relatively more noise than the other benchmarks.
+To reduce score noise, translation benchmarks were evaluated on the combined validation and test data splits, but there was still significant noise in the results.
+See figure 7 for individual benchmark visualizations.
 (a) ARC Challenge
@@
 Figure 5: Group I benchmark performance. Note the negative quadratic relationship with transformed PLW.
+The performance difference across different wp values for the two long-generation benchmarks was around twenty percentage points, in stark contrast to the less than two percentage point change for short-generation and multiple choice benchmarks.
+This suggests that PLW plays an important role in the ability to generate high quality text, and the optimal PLW for short-generation and long-generation benchmarks is clearly different.
+Also note that performance of LLaMA 2 models was in general higher than that of LLaMA 1 models and performance of AlpacaDataCleaned models were higher than that of AlpacaData models, validating the improvements of LLaMA 2 and AlpacaDataCleaned over their predecessors.
 (a) TruthfulQA-MC2
@@
 Figure 7: Group III benchmarks showed little relationship between performance and PLW.
-Appendix B Recommended Prompt Loss Weight
-In this section, we present additional figures and tables related to optimal prompt loss weights (PLWs) as predicted by fitted generalized additive models (GAMs).
-The optimal PLW predictions from the paper body were based on all benchmarks.
-Here we present three different relative aggregates: all benchmarks (labeled “All”), multiple choice benchmarks (labeled “MC”), and non-translation generation benchmarks (labeled “Gen”).
-We found that the translation benchmarks contributed little to the predictive power of the fitted GAMs.
-In an attempt to predict PLW values in a more relevant context, we did not include the translation benchmarks in the “Gen” model.
-However, they are still included in the “All” model to capture a wider range of tasks.
-See figures 7,8,9 for contour plots for the “All”, “MC”, and “Gen” benchmarks, respectively. The “All” contour plot is identical to the contour plot in the main body.
-See table 4 for a list of optimal PLWs over a range of completion-prompt ratios.
-Note that the predicted optimal PLWs confirm the conclusions from our regression analysis in section 5.2.
-Namely that PLW is important for short-completion data, but can be ignored for long-completion data.
+Appendix C Visualizations for Supplemental Experiments
+This appendix contains visualizations and additional details about the two supplemental experiments from section 6.
+(a) Simple Aggregate
+(b) Relative Aggregate
+(c) Relative Aggregate (Group I)
+(d) Relative Aggregate (Group II)
+Figure 8: Comparison of PLW with other regularization techniques (calculated for PLW $=0$ and PLW $=1$).
+(a) The simple aggregate.
+(b) The combined relative aggregate shows that models fine-tuned with fractional PLW on AlpacaDataShort outperformed models fine-tuned with alternative regularizations.
+(c) Fractional PLW performance is most extreme for multiple choice and short-generation benchmarks (group I).
+(d) Performance of fractional PLW models on group II benchmarks is less pronounced, with PLW-optimized models performing slightly worse than several other alternative metrics.
+C.1 Regularization Comparison
+For the first supplemental experiment, we wanted to investigate if PLW is necessary for fine-tuning on short-completion data or if another regularization technique could yield the same benefits.
+As explained in section 6.1, we chose to examine four types of regularization in addition to PLW, intentionally not evaluating KL divergence-based regularization due to the difficulty of applying it to LLM SIFT.
+Several aggregate scores for regularizations with example parameters are presented in figure 8, and best scores for each type of regularization is presented in table 4 in the main paper.
+Visualization of relative aggregate scores revealed that models fine-tuned with fractional PLW generated high scores on multiple-choice and short-generation benchmarks while long-generation benchmarks (AlpacaEval 1 and PandaLM) actually benefitted the most from alternative regularization methods.
+However, the effect on the multiple choice and short generation benchmarks was relatively strong, and the combined relative aggregate also showed maximal values for models fine-tuned with fractional PLW.
+Interestingly, most regularization methods performed better when coupled with PLW$=1$ than with PLW$=0$ except for label smoothing which performed marginally worse at PLW$=1$ than at PLW$=0$.
+C.2 Dataset Comparison
+(a) Original Datasets
+(b) Short Dataset Variants
+(c) UltraFeedbackBinarizedClean
+(d) UltraFeedbackBinarizedShort
+(e) DatabricksDolly
+(f) DatabricksDollyShort
+Figure 9: Relative aggregate scores for models fine-tuned on alternative instruction datasets.
+In the second supplemental experiment, we wanted to explore if the relationship between PLW and model performance on downstream tasks measured for AlpacaDataShort models existed for other fine-tuning datasets as well.
+See additional visualizations in figure 9.
+Appendix D Optimal Prompt Loss Weight
+(a) “All” Prediction
+(b) “MC” Prediction
+(c) “Gen” Prediction
+Figure 10: Best viewed digitally for improved resolution.
+In the main regression experiment, we showed that PLW is an important hyperparameter when fine-tuning on short-completion data but is effectively irrelevant when using long-completion data.
+In this appendix, we present several models for predicting an optimal PLW given a dataset’s $R_{g}$.
+These models are based on the AlpacaData dataset, AlpacaDataCleaned dataset, and several modified versions of AlpacaDataCleaned and therefore should be seen as an exercise rather than an authoritative reference on optimal PLW weights.
+We first repeated our SIFT experiments on two additional datasets: AlpacaDataMedium and AlpacaDataTiny to increase coverage of the parameter space.
+AlpacaDataMedium and AlpacaDataTiny have $R_{g}$ values of 1.0 and 0.042, respectively, and were generated using prompt inversion (see Appendix A) but selecting instances to modify in order to approach target $R_{g}$ values.
+We then fit several generalized additive model (GAMs) with a tensor smooth for the interaction between PLW and $R_{g}$.
+GAMs offer more flexible modeling than linear models but at the expense of interpretability.
+We fit our models using the R library mgcv and the following equation:
+score ~ te(w, r, k=3) + factor(b)”, where te is a full tensor product smooth, w is the untransformed PLW parameter, r is the $R_{g}$, k is the number of splines, and b is the benchmark task.
 $R_{g}$ Optimal PLW
@@
 2.5 1.000* 1.000* 1.000*
-2.0 0.239 1.000* 0.679*
-1.5 0.183 0.278 0.385*
-1.0 0.155 0.183 0.321*
-0.5 0.155 0.155 0.292*
+2.0 0.239 1.000* 0.679
+1.5 0.183 0.278 0.385
+1.0 0.155 0.183 0.321
+0.5 0.155 0.155 0.292
 Table 5: Optimal prompt loss weight (PLW) per completion-prompt length ratio $R_{g}$ on all benchmarks (“All”); multiple choice benchmarks (“MC”); and the combination of TruthfulQA-Gen, Alpaca Eval 1, and PandaLM benchmarks (“Gen”).
@@
 *The difference between the maximum and minimum predicted values at this ratio is less than 5% of the score range.
-Figure 7: “All” Ratio-PLW Interaction for all benchmarks.
-Figure 8: “MC” Ratio-PLW Interaction for multiple choice benchmarks.
-Figure 9: “Gen” Ratio-PLW Interaction for non-translation generation benchmarks.
-Appendix C Reproducibility
+Using the fitted w-r interaction, we then estimated optimal PLW value for a given completion-prompt ratio $R_{g}$. See figure 10(a) for a visualization of a GAM fitted on all benchmark tasks.
+Roughly, the fitted interaction term recommended using PLW $=0.155$ for small completion-prompt length ratios ($R_{g}\leq 1$) and up to PLW $=0.278$ for a $R_{g}=1.5$ for optimal performance across all tasks.
+This prediction is close to our regression-predicted value of 0.242.
+The interaction term also confirms our observations that PLW is less important for data with relatively long completions.
+Since the relationship between PLW and benchmark performance depends heavily on the type of benchmark task, we also fit GAMs for an aggregate of multiple choice benchmark scores (labeled “MC”) and generation benchmark scores (labeled “Gen”).
+We found that the translation benchmarks contributed little to the predictive power of the fitted GAMs and while their scores are included in the “All” GAM, we did not include them when fitting the “Gen” GAM.
+See figures 10(b) and 10(c) for contour plots for the “MC” and “Gen” benchmarks, respectively.
+Also see table 5 for a list of GAM-based optimal PLWs over a range of completion-prompt ratios. Again, predicted optimal PLWs confirmed the conclusions from our regression analysis in section 5.2.
+Appendix E Reproducibility
 This section provides technical details on all experiments and benchmarks for transparency and to encourage reproduction of results.
-To help with reproducibility, we will also upload the fine-tuned models, test generation outputs, and the AlpacaDataShort dataset to the HuggingFace Hub.
-C.1 Model Fine-Tuning
+To help with reproducibility, we will also upload the fine-tuned models, test generation outputs, and our modified datasets to xx.yy.zz.
+E.1 Model Fine-Tuning
 Model fine-tuning was performed with the Stanford Alpaca GitHub repository at www.github.com/tatsu-lab/stanford_alpaca/tree/761dc5b.
@@
 AlpacaDataCleaned can be found at www.github.com/gururise/AlpacaDataCleaned/tree/791174f and is labeled “alpaca_data_cleaned.json”.
-As noted above, we will release AlpacaDataShort in the near future to the HuggingFace Hub.
-C.2 Model Evaluation
+As noted above, we released AlpacaDataShort at xx.yy.zz.
+E.2 Model Evaluation
 We used three evaluation frameworks: EleutherAI’s Language Model Evaluation Harness (EEH), AlpacaEval 1, and PandaLM.
@@
 We modified the translation tasks at this commit to include an appropriate prompt to support zero-shot translation.
+These changes can be seen at xx.yy.zz.
 Though version 2 of AlpacaEval has recently been released, we used version 1 from the following commit https://github.com/tatsu-lab/alpaca_eval/tree/495b606.
@@
 For PandaLM, we used the commit at https://github.com/WeOpenML/PandaLM/tree/eb758c4 and evaluated on version 1 of the default test set (found at “data/testset-inference-v1.json” in the PandaLM repository).
-C.3 Regression
+E.3 Regression
 All statistical analysis and regression modeling was performed with R, version 4.3.0.
 We used the glmmTMB library, version 1.1.8, to perform generalized linear mixed modeling (GLMM) and validated results with the same library and with DHARMa, version 0.4.6.
-C.4 Causal Mechanism
+E.4 Causal Mechanism
 Most of the analysis performed to shed light on the causal mechanism should version and implementation agnostic.
@@
 ‘‘nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.4.0’’
-C.5 Predictive Model
-We fit several generalized additive models (GAMs) using the mgcv library, version 1.9-1 and the same version of R as above, version 4.3.0.
-Appendix D Artifact Licensing
+E.5 Supplemental Experiments
+The first supplemental experiment used the same experimental setup and data from the main experiment.
+Of the tested regularization methods, we used the weight decay implementation in PyTorch’s AdamW optimizer, attention dropout implemented by the LlamaAttention module from Transformers, and label smoothing supported by the Trainer class from Transformers.
+We manually implemented regularization based on the Minkowski distance by calculating the mean of the $p=1$ Minkowski distance between each pair of weight tensors from the PTLM and the trained model.
+The second experiment introduced two new datasets: UltraFeedbackBinarizedCleaned and DatabricksDolly.
+UltraFeedbackBinarizedCleaned can be found at https://huggingface.co/datasets/allenai/ultrafeedback_binarized_cleaned/tree/f304ce5.
+And DatabricksDolly can be found at https://huggingface.co/datasets/databricks/databricks-dolly-15k/tree/bdd27f4.
+E.6 Predictive Model
+We fit several generalized additive models (GAMs) in Appendix D using the mgcv library, version 1.9-1 and the same version of R as above, version 4.3.0.
+Appendix F Artifact Licensing
 We respected all licenses for artifacts and resources used in this research. Please see table 6 for an overview of primary resources and licenses.
@@
 Stanford Alpaca Apache 2.0 Model Training
+AlpacaDataCleaned Apache 2.0 Model Training
+Ultrafeedback Binarized Cleaned MIT Model Training
+databricks-dolly-15k CC BY-SA 3.0 Model Training
 LLaMA 1 LLaMA License Pre-trained model weights
 LLaMA 2 LLaMA 2 Community License Pre-trained model weights
+LLaMA 3 LLaMA 3 Community License Pre-trained model weights
 Mixtral 8x7B Apache 2.0 Model Evaluation

### v4

@@
 Instruction Fine-Tuning: Does Prompt Loss Matter?
-Mathew Huerta-Enochian  and Seung Yong Ko EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul {mathew, stephenko}@eq4all.co.kr
+Mathew Huerta-Enochian1thanks: Currently affiliated with the University of Oregon, research conducted while working at EQ4ALL.
+and Seung Yong Ko2 EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul 1mhuerate@uoregon.edu 2stephenko@eq4all.co.kr
 1 Introduction
@@
 While training, model parameters are updated by optimizing for next-token maximal likelihood classification.
-Most open sourced solutions for SFT either mask the prompt loss (for prefix language modeling) or use the entire sequence loss (for full language modeling) while some API providers support an explicit PLW parameter that allows users to apply fractional PLW during SIFT.
-The commonly-held notion is that fractional PLW helps stabilize when fine-tuning on data with short outputs.
+Most open sourced solutions for SIFT either mask the prompt loss (for prefix language modeling) or use the entire sequence loss (for full language modeling) while some API providers support an explicit PLW parameter that allows users to apply fractional PLW during SIFT.
+The commonly-held notion is that fractional PLW helps stabilize learning when fine-tuning on data with short outputs.
 Recently, however, OpenAI quietly removed support for their prompt_loss_weight parameter.
-The reason for the removal is unknown. Furthermore, to our knowledge, there has never been a proper study on the effects of PLW.
+The reason for the removal is unknown.
+Furthermore, to our knowledge, there has never been a proper study on the effects of PLW.
+[^1] Note that this research used a general implementation of PLW.
+OpenAI’s former prompt_loss_weight implementation could not be tested directly.
 We make the following contributions:
@@
 The goal of SIFT is to fine-tune a model to generate an appropriate completion for a given prompt.
-We define the generation ratio $R_{g}$ as the ratio of completion length to prompt length.
+We define the generation ratio $R_{g}$ as the ratio of completion length to prompt length (also referred to as the completion-prompt ratio).
 We then divide instruction data into two broad categories.
@@
 If $\lambda=0$, then PLW’s contribution to model performance is strictly negative, and if $\lambda=1$, then PLW contributes strictly positively to model performance.
-Note that $\lambda$ is not an intrinsic characteristic of the dataset, model architecture, or training algorithm.
+Note that $\lambda$ would not be an intrinsic characteristic of the dataset, model architecture, or training algorithm.
 Rather, it would depend on numerous factors and change for each task.
@@
 There are several qualitative performance trends that are of interest. For more thorough discussion and task-specific benchmark plots, see Appendix B.
-A group of four benchmarks (Arc Challenge, PIQA, TruthfulQA-Gen, and WinoGrande) clearly show the expected negative quadratic relationship between wp and performance of AlpacaDataShort models.
-Notably, AlpacaDataShort models outperform AlpacaData and AlpacaDataCleaned on this group given optimal PLW tuning.
-On the long-generation benchmarks, however, AlpacaDataShort models show a steadily-increasing trend.
-For these seven benchmarks and AlpacaDataShort models, using prompt loss masking (wp$=0$) almost always produced the worst scores.
-The wp-based performance increase was around twenty percentage points for both long-generation benchmarks.
-Contrasted with the less than two percentage point change for short-generation and multiple choice benchmarks, and it is evident that the optimal PLW for short-generation and long-generation benchmarks is different.
-Clear qualitative performance trends for the AlpacaData and AlpacaDataCleaned and for any model on the six translation benchmarks could not be identified.
+[…]
+Maximal wp-based performance increase was around twenty percentage points for both long-generation benchmarks and less than two percentage points for short-generation and multiple choice benchmarks.
+This difference in scale and the above difference in optimal PLW shows that the relationship between PLW and model performance is strongly dependent on the benchmark task.
+Clear qualitative performance trends for the AlpacaData- and AlpacaDataCleaned-trained models and for any model evaluated on the six translation benchmarks could not be identified.
 5.2 Regression
@@
 5.3 Causal Mechanism & Interpretation
+5.3.1 Training Stability
 To identify possible causal mechanisms, we first investigated the effects of PLW on training stability by analyzing training loss relative standard deviation (RSD) over five-step windows.
@@
 Considering that both AlpacaData and AlpacaDataCleaned models also showed a negative quadratic trend for mean loss RSD, we tentatively concluded that loss stability is not the driving factor for the modeled relationship.
+5.3.2 Weight Regularization
 We then checked if PLW was providing regularization to the weight update step, possibly improving performance by keeping weights close to the PTLM.
@@
 This is an interesting result, and we conclude that PTLM model weights were better preserved for small non-zero PLW despite high loss instability.
+5.3.3 Data Memorization
 We next explored how PLW affected training data memorization.
@@
 Corpus BLEU then decreased as PLW increased from 0.1.
-We also analyzed generation length on the AlpacaEval 1 test set which showed high deviations in generation length for the high loss RSD PLW values for LLaMA 2 only and a generally increasing length trend with PLW.
-Considering that AlpacaDataShort contains short completions, this strongly suggests that PLW contributes a regularizing effect and helps decrease overfitting.
-The positive relationship between performance and PLW on AlpacaEval 1 and PandaLM as opposed to the negative quadratic relationship seen on several of the other benchmarks can then be explained by a tradeoff between longer generations and decreased overfitting for high PLW and better preservation of PTLM weights for low PLW.
+We also analyzed generation length on the AlpacaEval 1 test set, which showed a generally increasing trend with PLW.222 Note that recent work Li et al.
+(2023); Dubois et al.
+(2024) has shown that AlpacaEval 1 has a preference for long generations, and we argue that improved AlpacaEval scores are not simply due to longer generations.
+First, while AlpacaEval performance showed a nearly strictly-increasing relationship with PLW, generation length did not.
+Second, we used Mixtral as the auto-evaluator which showed a much lower length preference than the default evaluator (0.63 and 0.75, respectively).
+Since AlpacaDataShort is dominated by short-completion instances, we concluded that non-zero PLW decreases overfitting by allowing the model to learn generation patterns from the prompt without negatively impacting instruction-completion alignment.
+5.3.4 Interpretation
+Based on the above analysis, we suggest that the causal mechanisms between PLW and downstream performance of models fine-tuned on short-completion data are
+  1. 1.
+    preservation of PTLM weights for small PLW and
+  2. 2.
+    reduced overfitting (and increased generation length) for large PLW.
+A tradeoff between these two mechanisms would explain the positive trend seen in the AlpacaEval and PandaLM benchmarks and the negative quadratic relationship in several of the other benchmarks.
 6 Supplemental Experiments
-In this section, we present supplemental PLW experiments that suggest that PLW cannot be replaced by alternative regularization techniques for SIFT and that the effects of PLW extend to other short-completion datasets.
+In this section, we present supplemental experiments that suggest that PLW cannot be replaced by alternative regularization techniques for SIFT and that the effects of PLW extend to other short-completion datasets.
 Since the translation benchmarks showed high levels of noise and unclear correlation with PLW in the main experiment, they were not included in supplemental experiments.
@@
 We tested weight decay, minimizing the Minkowski distance between PTLM weights and learned weights, dropout, and label smoothing.
-See Appendix C.1 for additional visualizations.
+See Appendix C.1 for visualizations.
 Note that we do not compare PLW with KL-divergence.
@@
 (b) UltraFeedbackShort and DatabricksDollyShort models.
-We repeated our main training experiment using two additional datasets, a cleaned and binarized version of UltraFeedback (Cui et al.
-(2023); denoted UltraFeedbackCleaned) and databricks-dolly-15k (Conover et al., 2023; denoted DatabricksDolly).
+We repeated our main training experiment using two additional datasets, a cleaned and binarized version of UltraFeedback (Cui et al., 2023; denoted UltraFeedbackCleaned) and databricks-dolly-15k (Conover et al., 2023; denoted DatabricksDolly).
 We also trained on prompt-inverted versions of both datasets, denoted UltraFeedbackShort and DatabricksDollyShort, respectively, and expanded analysis to include LLaMA 3 8B AI@Meta (2024) as a PTLM.
 See length statistics for these four datasets in table 1 and Appendix C.2 for additional visualizations.
-The combined relative aggregate scores for UltraFeedbackCleaned and DatabricksDolly and for UltraFeedbackShort and DatabricksDollyShort are visualized in figure 3.
+Note that compared with the unmodified datasets used in the main experiment, both UltraFeedback and DatabricksDolly have significantly lower generation ratios of $R_{g}=1.77$ and $R_{g}=0.83$, respectively.
+These datasets were chosen to demonstrate the effects of PLW when fine-tuning on data with relatively balanced generation ratios.
+The combined relative aggregate scores for models trained on UltraFeedbackCleaned and DatabricksDolly and for models trained on the prompt-inverted UltraFeedbackShort and DatabricksDollyShort datasets can be seen in figure 3.
 Visual inspection suggests that PLW affected learning for both groups of datasets.
-For the shortened variants, performance appears to have the same negative quadratic relationship with PLW as AlpacaDataShort models do in the main experiment.
-The relationship between PLW and performance for models fine-tuned on the original datasets is less clear but appears to have a generally increasing trend, unlike the long-completion datasets in the main experiment.
-We believe this is due to both UltraFeedbackCleaned and DatabricksDolly having much smaller completion-prompt length ratios $R_{g}$ than the AlpacaData and AlpacaDataCleaned datasets.
+For the shortened variants, performance appears to have the same negative quadratic relationship with PLW as the AlpacaDataShort models did in the main experiment.
+The relationship between PLW and performance for models fine-tuned on the unmodified datasets is weaker, but appears to be generally increasing.
+The clear relationship in the shortened data variants shows that the results of the main experiment extend to additional datasets.
+Furthermore, the weak relationship between PLW and performance of the unmodified data variants suggests that the effects of PLW are indeed dependent on the generation ratio.
+See Appendix D for an approach to predicting optimal PLW based on the generation ratio of the SIFT dataset.
 7 Conclusion
 In this study, we explored the effects of prompt loss weight (PLW) on LLM supervised instruction fine-tuning (SIFT).
-We found that PLW has a statistically significant effect on learning for our short-completion dataset, and proper tuning of PLW allowed short-completion-trained models to outperform long-completion-trained models on short-generation benchmarks.
+We found that PLW had a statistically significant effect on learning for our short-completion dataset, and proper tuning of PLW allowed short-completion-trained models to outperform long-completion-trained models on short-generation benchmarks.
 We showed that the causal mechanism was due to a balance between two different regularizing effects and not due to increased training stability as is commonly attributed.
@@
   2. 2.
-    Given the importance of PLW and given that many SIFT datasets and almost all natural language understanding (NLU) datasets are short-completion datasets, we warn SFT API providers about the need for a PLW parameter to adequately cover a full range of modeling applications.
+    Given the importance of PLW and given that many SIFT datasets and almost all natural language understanding (NLU) datasets are short-completion datasets, we warn SIFT API providers about the need for a PLW parameter to adequately cover a full range of modeling applications.
 Limitations
@@
     We analyzed prompt loss weighting (PLW) for instruction fine-tuning LLMs.
-We characterized three fine-tuning datasets by their relative completion-prompt length ratios and reported on the effect of PLW when training on each dataset.
+We characterized seven fine-tuning datasets by their relative completion-prompt length ratios and reported on the effect of PLW when training on each dataset.
 It would be helpful to extend this research to a wider range of datasets to increase the strength of our conclusions and create more complete guidelines for prompt loss weighting.
@@
   5. 5.
-    LLM-as-evaluator approaches like PandaLM and AlpacaEval 1 are still relatively new, and these approaches are being actively developed.
+    LLM-as-evaluator approaches like PandaLM and AlpacaEval are still relatively new, and these approaches are being actively developed.
 We chose to use Mixtral 8x7B as an auto-evaluator for AlpacaEval 1 due to budget limitations.
@@
 Acknowledgements
-This work was supported by the Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City (No.
+This work was supported by the Technology Innovation Program funded by the Korean Ministry of Trade, Energy, and Industry (MOTIE, Korea) (No.
+20014406, “Development of interactive sign language interpretation service based on artificial intelligence for the hearing impaired”) and the Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City (No.
 BA00000797, LLM-based sign language translation for weather forecasts).
-This research was possible thanks to the support of our coworkers at EQ4ALL.
 Appendices
@@
 In the first example above, prompt inversion changes the instance’s word-based completion-prompt ratio $R_{g}$ from $34/(7+0)=4.857$ to $7/(9+34)=0.163$.
-In order to experiment with short-completion fine-tuning datasets while also preserving a similar dataset to use as a reference, we propose using prompt inversion to rotate the instruction, input, and output fields of instances.
+In order to experiment with short-completion fine-tuning datasets, we propose using prompt inversion to rotate the instruction, input, and output fields of instances.
 Prompt inversion modifies an instance to use the original instruction as the new completion text and the original output as the new input text.
@@
 if there is an input field. See figure 4 for a visualization of this process.
-To generate short versions of our test instruction datasets, we use prompt inversion to modify every instance with a completion-prompt length ratio $R_{g}>1$, based on the tokenized lengths of each field and using prompt to mean the concatenation of the instruction and input fields as explained in section 2.1.
-Synthesis of the original input and output texts to predict the original instruction requires both language understanding and reasoning.
-It also does not train the model for long-generation (since new completions will generally be short) and thus does not prepare the model for train-of-thought generation that long-generation tasks intrinsically learn.
-Therefore, we propose that prompt-inverted instances are best suited for learning natural language understanding (NLU)-related tasks rather than general generation tasks.
+Synthesis of the original input and output texts to predict the original instruction requires both language understanding and reasoning, and prompt-inverted instances should be seen as natural language understanding (NLU) tasks.
+To generate short versions of the instruction datasets used in our experiments, we used prompt-inversion to modify every instance with a completion-prompt length ratio $R_{g}>1$, based on the tokenized lengths of each field (where “prompt” is the concatenation of the instruction and input fields as explained in section 2.1).
+Thus, given any long-completion dataset, a textually-similar short-completion dataset can be generated and used for comparison.
+Note that unless all instances have a generation ratio $R_{g}>1$, the resulting dataset will contain a mixture of unmodified instances and prompt-inverted instances.
 Appendix B Main Experiment Benchmarks
@@
 This section provides technical details on all experiments and benchmarks for transparency and to encourage reproduction of results.
-To help with reproducibility, we will also upload the fine-tuned models, test generation outputs, and our modified datasets to xx.yy.zz.
+To help with reproducibility, we will also uploaded the fine-tuned models, test generation outputs, and our modified datasets to the HuggingFace Hub and can be accessed at https://huggingface.co/collections/mathewhe/plw-66fe40fac6e586d8435bd563.
+Note that unless specified otherwise, default parameters were used for all training and testing.
 E.1 Model Fine-Tuning
-Model fine-tuning was performed with the Stanford Alpaca GitHub repository at www.github.com/tatsu-lab/stanford_alpaca/tree/761dc5b.
+Model fine-tuning was performed with the Stanford Alpaca GitHub repository at https://github.com/tatsu-lab/stanford_alpaca/tree/761dc5b.
 To experiment with prompt loss weight, we modified HuggingFace’s Transformers library to allow specifying a loss_weights parameter for LlamaForCausalLM’s forward method.
@@
 AlpacaData is available from the Stanford Alpaca repository.
-AlpacaDataCleaned can be found at www.github.com/gururise/AlpacaDataCleaned/tree/791174f and is labeled “alpaca_data_cleaned.json”.
-As noted above, we released AlpacaDataShort at xx.yy.zz.
+AlpacaDataCleaned can be found at https://github.com/gururise/AlpacaDataCleaned/tree/791174f and is labeled “alpaca_data_cleaned.json”.
+As noted above, AlpacaDataShort can be accessed at https://huggingface.co/collections/mathewhe/plw-66fe40fac6e586d8435bd563.
 E.2 Model Evaluation
 We used three evaluation frameworks: EleutherAI’s Language Model Evaluation Harness (EEH), AlpacaEval 1, and PandaLM.
-In an effort to match the current HuggingFace Open LLM leaderboard, we evaluated ARC Challenge, TruthfulQA-MC2, WinoGrande, and PIQA on the same EEH commit that the HuggingFace leaderboard uses: www.github.com/EleutherAI/lm-evaluation-harness/tree/b281b09 We also matched the number of shots with the number used for the HuggingFace leaderboard for ARC Challenge, TruthfulQA-MC2, and WinoGrande.
-TruthfulQA-Gen and all translation tasks were evaluated using a more recent commit at www.github.com/EleutherAI/lm-evaluation-harness/tree/b93c3bc.
+In an effort to match the current HuggingFace Open LLM leaderboard, we evaluated ARC Challenge, TruthfulQA-MC2, WinoGrande, and PIQA on the same EEH commit that the HuggingFace leaderboard uses: https://github.com/EleutherAI/lm-evaluation-harness/tree/b281b09 We also matched the number of shots with the number used for the HuggingFace leaderboard for ARC Challenge, TruthfulQA-MC2, and WinoGrande.
+TruthfulQA-Gen and all translation tasks were evaluated using a more recent commit at https://github.com/EleutherAI/lm-evaluation-harness/tree/b93c3bc.
 We modified the translation tasks at this commit to include an appropriate prompt to support zero-shot translation.
-These changes can be seen at xx.yy.zz.
+These changes can be seen at https://github.com/mathewhuen/plw_lm-evaluation-harness/compare/b93c3bc..1957d1a.
 Though version 2 of AlpacaEval has recently been released, we used version 1 from the following commit https://github.com/tatsu-lab/alpaca_eval/tree/495b606.
@@
 And DatabricksDolly can be found at https://huggingface.co/datasets/databricks/databricks-dolly-15k/tree/bdd27f4.
+The modified datasets UltraFeedbackShort and DatabricksDollyShort can be accessed at https://huggingface.co/collections/mathewhe/plw-66fe40fac6e586d8435bd563.
 E.6 Predictive Model